### v1

# Andreas Schlaginhaufen SYCAMORE, EPFL &Reda Ouhamma SYCAMORE, EPFL &Maryam Kamgarpour

Abstract
We study reinforcement learning from human feedback in general Markov decision processes, where agents learn from trajectory-level preference comparisons. A central challenge in this setting is to design algorithms that select informative preference queries to identify the underlying reward while ensuring theoretical guarantees. We propose a meta-algorithm based on randomized exploration, which avoids the computational challenges associated with optimistic approaches and remains tractable. We establish both regret and last-iterate guarantees under mild reinforcement learning oracle assumptions. To improve query complexity, we introduce and analyze an improved algorithm that collects batches of trajectory pairs and applies optimal experimental design to select informative comparison queries. The batch structure also enables parallelization of preference queries, which is relevant in practical deployment as feedback can be gathered concurrently. Empirical evaluation confirms that the proposed method is competitive with reward-based reinforcement learning while requiring a small number of preference queries.
1 Introduction
Reinforcement learning (RL) is a fundamental paradigm in machine learning, where agents learn to make sequential decisions by interacting with an environment to maximize cumulative rewards (Barto, 2021). RL has enabled advances in domains such as game play (Silver et al., 2017), robotics (Todorov et al., 2012), or autonomous driving (Lu et al., 2023). However, the practicality of RL is hindered by the challenge of designing rewards: crafting a reward function that aligns with human objectives is often difficult, and a misspecified reward function can lead to suboptimal or unsafe behavior (Amodei et al., 2016; Hadfield-Menell et al., 2017). This motivates the development of principled alternatives to manual reward design.
Rather than relying on manually specified reward functions, reinforcement learning from human feedback (RLHF) guides learning through preference feedback: at each step, a human oracle compares trajectories and indicates which is preferable (Christiano et al., 2017). This preference signal is often much easier to provide than engineering a reward function (Pereira et al., 2019; Lee et al., 2023). RLHF has proven to be effective in robotics (Jain et al., 2013) and, more recently, fine-tuning of large language models (Stiennon et al., 2020; Ziegler et al., 2019; Rafailov et al., 2023). This highlights the practical relevance of RLHF compared to reward-based learning.
Despite its empirical success, the theoretical foundations of RLHF are still in development. Existing works first studied the simpler setting of dueling bandits. In this context, the learner selects pairs of actions and observes noisy preference feedback (Yue et al., 2012; Komiyama et al., 2015). Classical algorithms for regret minimization in this setting include approaches based on zeroth-order optimization (Yue and Joachims, 2009) or the principle of optimism (Ailon et al., 2014). A key challenge in this setting is reducing the number of preference queries. For this purpose, several recent works propose strategic query selection strategies for dueling bandits (Das et al., 2024; Liu et al., 2024; Scheid et al., 2024; Mukherjee et al., 2024), often hinging on optimal experimental design mechanisms (Pukelsheim, 2006). However, such approaches are usually limited to finite-armed bandits, where the resulting optimization problems can be solved efficiently.
In the online RL setting, the theory of RLHF has received increasing attention, with several works establishing either regret or probably approximately correct (PAC) guarantees. PAC-RL methods aim to identify a near-optimal policy with high probability (Xu et al., 2020; Novoseller et al., 2020; Zhu et al., 2023), while regret-based approaches provide bounds on the cumulative reward during learning (Pacchiano et al., 2021; Chen et al., 2022; Wu and Sun, 2023). Similar to dueling bandits, a central challenge in RLHF is to actively select informative trajectory comparisons to drive learning. In RL, however, this active-learning problem presents additional difficulties: First, the learner cannot freely choose arbitrary state-action pairs or trajectories, but must reach them through exploration (Wagenmaker et al., 2022). Second, many existing approaches with guarantees (Pacchiano et al., 2021; Zhan et al., 2024) rely on maximizing an exploration bonus involving a norm of state distributions – a problem which is known to be computationally intractable even in tabular settings (Efroni et al., 2021).
To sidestep these challenges, another line of work focuses on RLHF with offline data. In this setting, learning proceeds over a fixed pre-collected dataset of trajectory preferences (Zhu et al., 2023; Zhan et al., 2023). Although this offline paradigm avoids the need for online exploration and active query selection, it depends critically on having access to sufficiently diverse and informative preference data a priori (Rashidinejad et al., 2021; Xie et al., 2021; Zanette et al., 2021; Zanette, 2023) — a requirement that can be difficult to meet in practice. Hence, this merely shifts the exploration and active learning challenges to the data collection phase.
Despite progress on statistical guarantees in RLHF, two key challenges remain open: the design of tractable algorithms for active preference query selection and reducing the workload of human preference annotators. In existing approaches, a human must provide feedback at every round, which is impractical in real-world applications. Our goal in this work is to develop RLHF algorithms that are computationally efficient, reduce the demand for human feedback, and actively select informative queries. Some recent work has made progress on tractability. For instance, Wu and Sun (2023) proposes a randomized exploration algorithm with regret guarantees, but their method is limited to linear dynamics. In parallel, Wang et al. (2023) introduces a general reduction from RLHF to standard RL and establishes PAC-style guarantees under RL oracle access. However, neither approach addresses the open challenges of reducing feedback requirements or enabling active query selection.
Contributions
In this work, we focus on reinforcement learning from human feedback (RLHF) and develop meta-algorithms that reduce the RLHF problem to standard RL by leveraging existing RL algorithms as subroutines. Our approach combines randomized exploration for tractability, lazy updates to reduce human workload, and experimental design to actively select informative preference queries. We provide both regret and PAC-style guarantees under RL oracle assumptions. Our contributions are as follows:
  • We propose general meta-algorithms for RLHF using RL oracles, and provide provable regret and PAC guarantees in general MDPs.
  • We present a second meta-algorithm with better scalability and query efficiency thanks to: Lazy updates, inspired by linear bandits (Abbasi-Yadkori et al., 2011), which enables parallelization of the preference oracle calls; Greedy optimal design, which selects high-quality preference queries and improves sample efficiency.
  • We provide empirical results showing that: Our algorithm is implementable and competitive with reward-based RL; The improved algorithm achieves comparable performance while significantly reducing the query complexity. [^1] The code is openly accessible at https://github.com/andrschl/isaac_rlhf.
2 Preliminaries
Notation
[…]
Setting
We consider an infinite-horizon [^2] Our results extend directly to the finite-horizon setting as well. However, we focus on the infinite-horizon discounted setting, as it is more commonly encountered in deep reinforcement learning. Markov decision process (MDP) =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\nu_{0},P,r,\gamma\}$ with compact state and action spaces $\mathcal{S}\subset\mathbb{R}^{n}$ and $\mathcal{A}\subset\mathbb{R}^{m}$, respectively, initial state distribution $s_{0}\sim\nu_{0}$, transition law $s_{h+1}\sim P(\cdot|s_{h},a_{h})$, and discount rate $\gamma\in(0,1)$. We assume a linear reward model $r_{\theta^{*}}(s,a):=\langle\theta^{*},\phi(s,a)\rangle$, where $\left\lVert\theta^{*}\right\rVert\leq B$ and $\phi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$ is a feature mapping such that $\max_{s,a}\left\lVert\phi(s,a)\right\rVert\leq L$. We denote the set of all trajectories as $\mathcal{T}:={(\mathcal{S}\times\mathcal{A})}^{\infty}$, and the distribution over $\mathcal{T}$ induced by a stationary Markov policy $\pi:\mathcal{S}\to\Delta_{\mathcal{A}}$ as $\mathbb{P}_{\pi}$. For a trajectory $\tau=(s_{0},a_{0},s_{1}\ldots)\in\mathcal{T}$, we denote the discounted sum of features by $\phi(\tau):=\sum_{h=0}^{\infty}\gamma^{h}\phi(s_{h},a_{h})$ and the feature expectation of a policy $\pi$ by $\phi(\pi):=\mathbb{E}_{\tau\sim\mathbb{P}_{\pi}}\left[{\phi(\tau)}\right]$. Furthermore, given a reward parameter $\theta$, we denote the value of a policy $\pi$ by $V_{\theta}^{\pi}:=\mathbb{E}_{\tau\sim\mathbb{P}_{\pi}}\left[{\sum_{h=0}^{\infty}\gamma^{h}r_{\theta}(s_{h},a_{h})}\right]=\langle\theta,\phi(\pi)\rangle$ and the optimal value by $V_{\theta}^{*}:=\max_{\pi}V_{\theta}^{\pi}$.
Interaction protocol
For each round $t=1,\ldots,T$ of RLHF, a learner, the MDP, and a preference oracle interact as follows. The learner selects two policies $\pi_{t}$ and $\pi^{\prime}_{t}$, and executes them to obtain two trajectories $\tau_{t}\sim\mathbb{P}_{\pi_{t}}$ and $\tau_{t}^{\prime}\sim\mathbb{P}_{\pi_{t}^{\prime}}$. Subsequently, the learner may query the preference oracle, which returns a binary label $y_{t}=\mathbbm{1}\left({\tau_{t}\succ\tau^{\prime}_{t}}\right)\in\left\{{0,1}\right\}$. The label equals one if the trajectory $\tau_{t}$ is preferred over $\tau^{\prime}_{t}$, denoted as $\tau_{t}\succ\tau^{\prime}_{t}$, and zero otherwise. Each such interaction is one RLHF round.
Preference model
We consider a stochastic preference model characterized by a preference function $\mathcal{P}:\mathcal{T}\times\mathcal{T}\to[0,1]$, assigning to each pair of trajectories $\tau,\tau^{\prime}\in\mathcal{T}$ the probability $\mathcal{P}(\tau\succ\tau^{\prime})$ of preferring $\tau$ to $\tau^{\prime}$. We make the following assumption about the preference model.
Assumption 2.1 (Bradley–Terry model).
The preference function $\mathcal{P}$ satisfies for all $\tau,\tau^{\prime}\in\mathcal{T}$
$\mathcal{P}(\tau\succ\tau^{\prime})=\sigma\left({\langle\theta^{*},\phi(\tau)-\phi(\tau^{\prime})\rangle}\right),$ (1)
where $\sigma(x)=1/(1+e^{-x})$ denotes the sigmoid function.
This preference model is a special case of the Plackett-Luce model (Plackett, 1975; Luce et al., 1959), and is commonly used in the dueling bandit setting as well as the RLHF framework (Yue et al., 2012; Christiano et al., 2017; Ouyang et al., 2022). In particular, this model satisfies the strong stochastic transitivity property (Strzalecki, 2025). That is, for any three trajectories $\tau_{1}\succ\tau_{2}\succ\tau_{3}$ we have $\mathcal{P}(\tau_{1}\succ\tau_{3})>\max\{\mathcal{P}(\tau_{1}\succ\tau_{2}),\mathcal{P}(\tau_{2}\succ\tau_{3})\}$, capturing the monotonicity of preferences.
Remark 2.2.
In practice, we cannot compare trajectories of infinite length. Fortunately, many environments terminate in finite time, and otherwise one may truncate each trajectory at horizon $H=\mathcal{O}(\log_{\gamma}(\varepsilon))$, introducing at most an $\varepsilon$ error in value estimates (see e.g. (Schlaginhaufen and Kamgarpour, 2024)). For simplicity, however, we omit this truncation step in our presentation.
Regret
To assess the learner’s online performance, we consider the cumulative regret
$R(T)=\sum_{t=1}^{T}\frac{\left({V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}}}\right)+(V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi^{\prime}_{t}})}{2}=\frac{1}{2}\sum_{t=1}^{T}\left({2V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}}-V_{\theta^{*}}^{\pi_{t}^{\prime}}}\right).$
Cumulative regret has been widely adopted in the RL and RLHF literature (Abbasi-Yadkori et al., 2011; Zanette et al., 2020; Wang et al., 2023; Zhan et al., 2024). However, cumulative regret doesn’t provide us with a guarantee of the last iterate’s suboptimality. As a second metric, we therefore also consider the suboptimality of an output policy.
Suboptimality
The suboptimality of an output policy $\hat{\pi}$ is defined by
$\mathsf{SubOpt}(\hat{\pi}):=V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\hat{\pi}}.$
Suboptimality has previously been considered as a performance metric for offline RLHF Zhu et al. (2023), contextual bandits Das et al. (2024), and online RLHF (Wang et al., 2023). In the following, we propose a meta-algorithm that features two variants: one with theoretical guarantees on cumulative regret, and another specifically ensuring a bound on last-iterate suboptimality.
3 Randomized Preference Optimization
3.1 Algorithm
Our algorithm proceeds in three essential steps. First, the preference feedback data is used to estimate the reward parameter using maximum likelihood estimation. Then, a reward parameter is sampled from a Gaussian distribution, which is reminiscent of linear Thompson sampling (Abeille and Lazaric, 2017). Finally, an RL oracle is used to find an approximately optimal policy for the sampled reward parameter, and we query the preference oracle by comparing one trajectory from this new policy against one from the previous policy.
Maximum likelihood estimation
Considering our preference model (1), a standard approach for estimating the reward parameter $\theta^{*}$ is via maximum likelihood estimation. Given a pair of trajectories $\tau_{k}=(s_{h,k},a_{h,k})_{h=0}^{\infty}$ and $\tau_{k}=(s^{\prime}_{h,k},a^{\prime}_{h,k})_{h=0}^{\infty}$ we consider the design points $x_{k}:=\phi(\tau_{k})-\phi(\tau^{\prime}_{k})=\sum_{h=0}^{\infty}\gamma^{h}(\phi(s_{h,k},a_{h,k})-\phi(s^{\prime}_{h,k},a^{\prime}_{h,k}))$ and the preference labels $y_{k}=\mathbbm{1}(\tau_{k}\succ\tau^{\prime}_{k})$. In round $t$, the preference dataset is $\mathcal{D}_{t}=\{(x_{k},y_{k})\}_{k=1}^{t-1}$ and the corresponding (constrained) maximum likelihood estimator (MLE) is given by $\hat{\theta}_{t}=\operatorname*{arg\,min}_{\left\lVert\theta\right\rVert\leq B}\mathcal{L}_{\mathcal{D}_{t}}(\theta)$, where
$\mathcal{L}_{\mathcal{D}_{t}}(\theta):=-\sum_{(x,y)\in\mathcal{D}_{t}}\left[{y\log\sigma(\langle\theta,x\rangle)+(1-y)\log\sigma(-\langle\theta,x\rangle)}\right],$ (2)
is the negative log-likelihood of the Bradley-Terry model (1). The loss function (2) is the familiar logistic loss from logistic regression (Shalev-Shwartz and Ben-David, 2014). In particular, it is a convex problem that can be solved efficiently using standard methods such as LBFGS (Liu and Nocedal, 1989). Moreover, we have the following time-uniform confidence result.
Lemma 3.1.
Let $\lambda>0$ and define the design matrix at time $t$ given by $V_{t}=\lambda I+\sum_{s=1}^{t-1}x_{t}x_{t}^{\top}$. Then, with probability $1-\delta$, for all $t\geq 1$, the true reward parameter $\theta^{*}$ is contained in the ellipsoid
$\mathcal{E}_{t}(\delta):=\left\{{\theta:\left\lVert\theta-\hat{\theta}_{t}\right\rVert_{V_{t}}^{2}\leq\beta_{t}(\delta)^{2}:=\mathcal{O}\left({\kappa\left[{\log\left({\frac{1}{\delta}}\right)+d\log\left({\frac{t}{d}}\right)}\right]+\lambda}\right)}\right\}.$
Here, $\kappa:=\max_{\theta\in\mathcal{B}^{d}(B),x\in\mathcal{B}^{d}(2LH_{\gamma})}1/\dot{\sigma}\left({\langle\theta,x\rangle}\right)$ denotes the Lipschitz constant of the inverse sigmoid function, and $H_{\gamma}=(1-\gamma)^{-1}$ the effective horizon of the MDP.
The above lemma hinges on a likelihood-ratio confidence set from Lee et al. (2024). The proof and the precise constants are deferred to Appendix A.
Remark 3.2.
Compared to the standard analysis of stochastic linear bandits Abbasi-Yadkori et al. (2011), our parameter $\beta_{t}$ includes an additional factor of $\sqrt{\kappa}$, which arises naturally due to preference-based feedback. This result improves upon the bound provided by Zhu et al. (2023), which incurs a larger factor of $\kappa$ instead of $\sqrt{\kappa}$. While the $\sqrt{\kappa}$ factor can theoretically be avoided by constructing confidence sets using the Hessian of the negative log-likelihood $\mathcal{L}_{\mathcal{D}_{t}}$ (Lee et al., 2024; Das et al., 2024), it reappears in the regret bounds as shown in (Das et al., 2024). Thus, we adopt confidence sets based on $V_{t}$, as both $V_{t}$ and its inverse can be efficiently updated via rank-one operations, unlike Hessian-based approaches.
Randomized exploration
Many approaches to regret minimization and pure exploration in RLHF often rely on maximizing an exploration bonus $\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}$. Although such methods yield provable guarantees for regret (Pacchiano et al., 2021) or last-iterate suboptimality (Das et al., 2024), they are computationally intractable in RL settings (see Appendix E.2 for a discussion). To address this, we adopt a randomized exploration scheme inspired by Thompson sampling algorithms for linear bandits. In line with Abeille and Lazaric (2017), we sample the reward parameter from an inflated version of the confidence set defined in Lemma 3.1, which produces a computationally efficient alternative to optimism-based approaches. Furthermore, we also show that this randomized strategy extends to the pure exploration setting.
RL oracle
With the objective of a meta-algorithm, we assume access to the following RL oracle.
Assumption 3.3 (PAC-RL oracle).
We assume access to an $(\varepsilon,\delta)$-PAC oracle, $\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}$, for the RL problem. That is, a polynomial-time algorithm that produces for every $\varepsilon>0$, $\delta>0$, and $\theta\in\mathbb{R}^{d}$ a policy $\pi=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}(\theta,\varepsilon,\delta)$ such that with probability at least $1-\delta$ we have $V^{*}_{\theta}-V_{\theta}^{\pi}\leq\varepsilon$.
This assumption of a PAC-RL oracle is satisfied in several settings, including tabular and linear MDPs (Dann et al., 2019; Ménard et al., 2021; Al Marjani et al., 2021; Wagenmaker et al., 2022; He et al., 2021). Moreover, it is relatively mild compared to stronger oracles considered in the RLHF literature (Zhan et al., 2024), such as reward-free algorithms (Wang et al., 2020; Kaufmann et al., 2021; Ménard et al., 2021). In practice, common choices for $\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}$ are policy optimization methods such as proximal policy optimization (PPO) (Schulman et al., 2017) or soft actor critic (Haarnoja et al., 2018), which have shown strong empirical performance in continuous control and large-scale applications such as training large language models.
Algorithm statement
Our algorithm RPO presented below comes in two variants: a variant that balances exploration and exploitation for regret minimization ($\alpha=1$), and a variant for pure exploration ($\alpha=0$). In line 4, we sample a reward parameter from a confidence set inflated by $\sqrt{d}$. Then, in line 5, we compute the policy $\pi_{t}$ using an RL oracle. Finally, in line 10, we use the new preference data to update the reward parameter using maximum likelihood.
Input: Number iterations $T$, confidence $\delta>0$, regularization $\lambda>0$, algorithm type $\alpha\in\left\{{0,1}\right\}$.
1 Set $\varepsilon=1/\sqrt{T}$ and $\delta^{\prime}=\delta/5$.
2 Initialize design matrix $V_{1}=\lambda I$, preference data set $\mathcal{D}_{1}=\emptyset$, parameter $\hat{\theta}_{1}$, and policy $\pi_{0}$.
3 for $t=1,2,\dots,T$ do
       $\tilde{\theta}_{t}\sim\mathcal{N}\big{(}\alpha\hat{\theta}_{t},\beta_{t}(\delta^{\prime})^{2}V_{t}^{-1}\big{)}$;
        // Reward sampling
       $\pi_{t}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\big{(}\tilde{\theta}_{t},\varepsilon,\delta^{\prime}/T\big{)},\pi^{\prime}_{t}=\pi_{t-1}$;
        // Update policy with RL
4       $x_{t}=\phi(\tau_{t})-\phi(\tau_{t}^{\prime})$ with $\tau_{t}\sim\mathbb{P}_{\pi_{t}},\tau^{\prime}_{t}\sim\mathbb{P}_{\pi^{\prime}_{t}}$;
       $y_{t}=\mathbbm{1}(\tau_{t}\succ\tau^{\prime}_{t})$;
        // Preference feedback
5       $\mathcal{D}_{t+1}=\mathcal{D}_{t}\cup\left({y_{t},x_{t}}\right)$;
6       $V_{t+1}=V_{t}+x_{t}x_{t}^{\top}$;
       $\hat{\theta}_{t+1}\in\operatorname*{arg\,min}_{\left\lVert\theta\right\rVert\leq B}\mathcal{L}_{\mathcal{D}_{t+1}}(\theta)$ ;
        // Reward estimation
7
8 end for
Output: Policy $\hat{\pi}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\bar{\theta}_{T},\varepsilon,\delta^{\prime}}\right)$ with $\bar{\theta}_{T}=\frac{1}{T}\sum_{t=1}^{T}\hat{\theta}_{t}$.
Algorithm 1 Randomized Preference Optimization (RPO)
3.2 Theoretical results
We analyze the regret of Algorithm 1 and the suboptimality of its output policy $\hat{\pi}$.
3.2.1 Regret analysis
We show that Algorithm 1 with $\alpha=1$ has sublinear regret with high probability.
Theorem 3.4.
Using Algorithm 1 with $\alpha=1$, it holds for any $\delta>0$ and $T\in\mathbb{N}$, with probability at least $1-\delta$ that
$R(T)\leq\mathcal{O}\left(\sqrt{\kappa d^{3}T\log(dT/\delta)^{3}}\right).$
The regret bound of Theorem 3.4 matches the best existing bounds for algorithms with randomized exploration in reinforcement learning, see (Efroni et al., 2021; Ouhamma et al., 2023). In addition, due to learning from preferences, we have an extra $\sqrt{\kappa}$ factor (see Remark 3.2), which is in line with other recent work on RLHF (Wu and Sun, 2023; Das et al., 2024).
Comparison with prior work
For episodic tabular MDPs Pacchiano et al. (2021) prove a regret bound of $\widetilde{\mathcal{O}}(\kappa d\sqrt{T})$. Similarly, Chen et al. (2022) considers episodic linear MDPs and derives a regret bound of $\widetilde{\mathcal{O}}(d\sqrt{HT})$, avoiding dependence on $\kappa$ by assuming a linear preference model. However, these approaches rely on a type of optimism which is computationally intractable (see Appendix E.2). Similar to us, Wu and Sun (2023) avoids this challenge by resorting to randomized exploration and proves a $\widetilde{\mathcal{O}}(d^{3}\sqrt{\kappa T})$ regret bound for linear MDPs. Compared to Wu and Sun (2023), our analysis improves the dependence on dimensionality from $d^{3}$ to $d^{3/2}$ and avoids the need for truncation techniques on the value function. Furthermore, our settings differ in two key points: First, we assume access to an RL oracle without restricting the class of MDPs, whereas Wu and Sun (2023) considers linear MDPs. Second, their approach is model-based, while Algorithm 1 is oracle-based and can accommodate both model-based and model-free implementations.
Proof idea
The full proof of this Theorem is provided in Appendix B. Analogous to the analysis of linear Thompson sampling (Abeille and Lazaric, 2017), the main idea is to control the regret by showing that randomized exploration ensures a constant probability of optimism. However, compared to the linear bandit analysis, our setting comes with additional challenges: First, due to preference-based learning we require a different regret decomposition accounting for the reference policy. Second, as we observe preference feedback on trajectories rather than policies, we need to apply Freedman’s inequality (see Lemma A.9) to control the deviation between expected and observed features. Lastly, as we assume a PAC RL oracle – in place of an exact maximization oracle – we need to carefully track the resulting approximation error.
3.2.2 Suboptimality gap
The pure exploration version of RPO with $\alpha=0$, in contrast to $\alpha=1$, may incur linear regret during learning as the policy $\pi_{t}$ can be highly suboptimal. However, Theorem 3.5 below shows that it outputs a policy that is $\tilde{\mathcal{O}}(1/\sqrt{T})$-optimal.
Theorem 3.5.
Using Algorithm 1 with $\alpha=0$, it holds for any $\delta>0$ and $T\in\mathbb{N}$, with probability at least $1-\delta$ that
$\mathsf{SubOpt}\left({\hat{\pi}}\right)=\mathcal{O}\left({\sqrt{\frac{\kappa d^{3}}{T}\log\left({\frac{dT}{\delta}}\right)^{3}}}\right).$
In other words, we need $\tilde{\mathcal{O}}(\kappa d^{3}/\varepsilon^{2})$ iterations to output an $\varepsilon$-optimal policy with high probability.
Except for the extra $\sqrt{d}$ dependency, which is inherent to approaches based on Thompson sampling, [^3] Recently, Abeille et al. (2025) show that for certain classes of linear bandits the additional $\sqrt{d}$ factor can be avoided. However, their assumptions are not directly applicable to our preference-based RL setting. we match the last iterate guarantee proposed by Das et al. (2024) for contextual linear bandits, but with an algorithm that is tractable in the full RL setting.
Comparison with prior work
Few works provide suboptimality guarantees in preference-based RL. Wang et al. (2023) propose a meta-algorithm interfacing with a PAC-RL oracle that outputs an $\varepsilon$-optimal policy after $\widetilde{\mathcal{O}}(\kappa^{2}d^{3}/\varepsilon^{2})$ queries. A different approach by Zhan et al. (2024) leverages optimal design to prove a bound of $\widetilde{\mathcal{O}}((|S|^{2}|A|d+\kappa^{2}d^{2})/\varepsilon^{2})$, but their method relies on an intractable maximization oracle. In comparison, our algorithm achieves a bound of $\widetilde{\mathcal{O}}(\kappa d^{3}/\varepsilon^{2})$, improving the dependence on $\kappa$ over both prior results. The additional factor of $d$ compared to Zhan et al. (2024) is expected, as our method is randomized rather than optimistic, see (Abeille and Lazaric, 2017). Notably, to the best of our knowledge, this is the first algorithm, whether in preference-based or reward-based RL, that provides last-iterate guarantees using randomized exploration.
Proof idea
The proof of Theorem 3.5, presented in Appendix C, builds on Das et al. (2024)’s suboptimality analysis for the contextual bandits setting. However, to sidestep the intractability of maximizing an exploration bonus over policies, we leverage randomized exploration (Abeille and Lazaric, 2017) to ensure a constant probability of optimism. This allows us to derive a bound on the output policy’s suboptimality that mirrors the regret bound, without needing additional assumptions.
3.3 Practical limitations
While Algorithm 1 is tractable and statistically efficient, it presents certain limitations. First, invoking an RL oracle at every round (line 5) can cause high latency, especially in continuous state spaces. Second, issuing preference queries at each round (line 7) is impractical when the oracle is a human, due to the need for continuous feedback. Finally, requesting a label for all trajectory pairs can be expensive and inefficient, as many comparisons are uninformative.
The next section presents a refined algorithm addressing these limitations. The new approach decouples trajectory collection from query selection, and queries only the most informative comparisons.
4 A practical algorithm with efficient query selection
We present Algorithm 2, an improved method for preference collection and active query selection.
4.1 Algorithm
As discussed earlier, we design Algorithm 2 by using lazy updates to collect a batch of trajectory pairs, then applying optimal design to select the informative queries from the batch.
Input: Number iterations $T$, confidence $\delta>0$, regularization $\lambda>1$, constant $C>0$.
1 Set $\varepsilon=1/\sqrt{T}$, $\delta^{\prime}=\delta/5$, $t_{\operatorname{stop}}=1$.
2 Initialize design matrices $V_{0}=W_{0}=\lambda I$, preference datasets $\mathcal{D}=\mathcal{D}_{0}=\emptyset$, parameter $\hat{\theta}_{0}$.
3 for $t=1,2,\dots,T$ do
4       if $\det(W_{t})>(1+C)\det(W_{t_{\operatorname{stop}}})$ then
             $\mathcal{D}_{\text{opt}},V_{t}=\operatorname{D-OptDes}(\mathcal{D},V_{t_{\operatorname{stop}}},\det(W_{t}))$ ;
              // Greedy D-Optimal design
5             $\mathcal{D}_{t}=\mathcal{D}_{t_{\operatorname{stop}}}$;
6             for $(\tau,\tau^{\prime})\in\mathcal{D}_{\text{opt}}$ do
                   $x=\phi(\tau)-\phi(\tau^{\prime}),y=\mathbbm{1}(\tau\succ\tau^{\prime})$ ;
                    // Preference feedback
7                   $\mathcal{D}_{t}=\mathcal{D}_{t}\cup\{(x,y)\}$;
8             end for
            $\hat{\theta}_{t}\in\operatorname*{arg\,min}_{\left\lVert\theta\right\rVert\leq B}\mathcal{L}_{\mathcal{D}_{t}}(\theta)$ ;
              // Reward estimation
9             $t_{\operatorname{stop}}=t$, $\mathcal{D}=\emptyset$, $\pi^{\prime}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\hat{\theta}_{t},\varepsilon,\delta^{\prime}/T}\right)$;
10
11       end if
      $\tilde{\theta}_{t}\sim\mathcal{N}(\hat{\theta}_{t_{\operatorname{stop}}},\beta_{t_{\operatorname{stop}}}(\delta^{\prime})^{2}V_{t_{\operatorname{stop}}}^{-1})$;
        // Reward sampling
       $\pi_{t}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\tilde{\theta}_{t},\varepsilon,\delta^{\prime}/T}\right)$;
        // Update policy with RL
12       $x_{t}=\phi(\tau_{t})-\phi(\tau_{t}^{\prime})$ with $\tau_{t}\sim\mathbb{P}_{\pi_{t}},\tau^{\prime}_{t}\sim\mathbb{P}_{\pi^{\prime}}$;
13       If $t=t_{\operatorname{stop}}$, then $W_{t+1}=V_{t}+x_{t}x_{t}^{\top}$, else: $W_{t+1}=W_{t}+x_{t}x_{t}^{\top}$;
14       $\mathcal{D}=\mathcal{D}\cup\{(\tau_{t},\tau^{\prime}_{t})\}$;
15
16 end for
Algorithm 2 Lazy Randomized Preference Optimization with Optimal Design (LRPO-OD)
Lazy updates
We use an idea from Abbasi-Yadkori et al. (2011) to collect many trajectory pairs without querying the preference oracle. The modification compared to Algorithm 1 is collecting trajectories without updating the MLE $\hat{\theta}_{t_{\operatorname{stop}}}$ until the information gain, represented by $\det(V_{t})$, increases by a multiplicative constant; see line 4 of Algorithm 2. We show that this procedure limits the number of batches to $\mathcal{O}(\log(T))$. In other words, the average (over batches) size of a given batch is of order $\mathcal{O}(T/\log(T))$. A key advantage of this lazy update structure is that the preference queries (line 8) can be collected in parallel across all trajectory pairs within a batch. This significantly reduces the workload of the preference oracle, e.g. a human annotator, by eliminating the need for round-by-round feedback. In particular, the algorithm no longer pauses at each timestep to wait for preference labels. To our knowledge, this is the first approach to support parallelization of queries in online preference-based reinforcement learning, even in settings as simple as dueling bandits.
D-Optimal design
To select informative preference queries from the collected trajectories above, we leverage tools from optimal experimental design. Specifically, we apply an approximate D-optimal design criterion to each collected batch of trajectory pairs; see Appendix E.1 for background on D-optimal design. Given the current matrix $V_{t_{\operatorname{stop}}}$ and the set of candidate trajectory pairs $\mathcal{D}$, we use a greedy algorithm to solve the following maximization problem:
$\max_{\{n_{x}\}}\log\det\left({V_{t_{\operatorname{stop}}}+\sum_{x\in\mathcal{D}}n_{x}xx^{\top}}\right)\quad\text{ subject to }\sum_{x\in\mathcal{D}}n_{x}=|\mathcal{D}|,\>\>n_{x}\in\mathbb{N}.$
Due to the submodularity of the $\log\det$ function for $\lambda$ greater than one, the greedy procedure of Algorithm 3 achieves an $(1-1/e)$-approximation to the optimal solution; see (Nemhauser et al., 1978; Krause et al., 2008) and Appendix E.1.
Another key feature of Algorithm 3 is that the while loop is stopped early if $\det(V)$ exceeds the threshold value. This threshold is set as the determinant of the naive design, where every trajectory pair is queried once. When this early termination is satisfied, Algorithm 2 requires fewer samples than Algorithm 1. In particular, since the optimal design maximizes the information gain (measured by $\det(V)$), this termination condition is expected to be satisfied frequently in practice.
Input: Dataset $\mathcal{D}$, current design matrix $V$, threshold for the determinant value $\alpha$.
1 Initialize and dataset $\mathcal{D}_{\text{opt}}=\emptyset$
2 while $\det(V)<\alpha$ and $|\mathcal{D}_{\text{opt}}|\leq|\mathcal{D}|$ do
3       $(\tau,\tau^{\prime})=\operatorname*{arg\,max}_{(\tau,\tau^{\prime})\in\mathcal{D}}\det(V+(\phi(\tau)-\phi(\tau^{\prime}))(\phi(\tau)-\phi(\tau^{\prime}))^{\top})$
4       $V=V+xx^{\top},\text{ where }x=\phi(\tau)-\phi(\tau^{\prime});\>\mathcal{D}_{\text{opt}}=\mathcal{D}_{\text{opt}}\cup\{(\tau,\tau^{\prime})\}$
5 end while
Output: Curated dataset $\mathcal{D}_{\text{opt}}$, design matrix $V$.
Algorithm 3 Greedy D-Optimal Design
4.2 Theoretical result
We now provide our high probability regret bound for Algorithm 2.
Theorem 4.1.
Instantiating Algorithm 2 with $C>0$ and $\lambda>1$, it holds for any $\delta>0$ and $T\in\mathbb{N}$, with probability at least $1-\delta$ that
$R(T)\leq\mathcal{O}\left(\sqrt{(1+C)\kappa d^{3}T\log(dT/\delta)^{3}}\right).$
In addition, the number of times the condition of line 4 holds is at most $\frac{d}{\log(1+C)}\log\left({1+\frac{T(LH_{\gamma})^{2}}{\lambda d}}\right)$. Therefore, the size of the batches is on average of order $\tilde{\mathcal{O}}(T/(d\log(T)))$.
Compared to Theorem 3.4, the above regret bound increases only by constant factors. Regarding the number of preference queries, Sekhari et al. (2023) shows that at least $\Omega(T)$ queries are required to achieve $\mathcal{O}(\sqrt{T})$ regret in the worst case. However, optimal design may lead to significantly fewer queries in favorable instances, as demonstrated in our experiments. Importantly, approximate optimal design does not compromise our guarantees, as it ensures a $1/(1-1/e)$ approximation to the optimal.
Proof idea
We present here the key ideas of the proof of Theorem 4.1, the full proof is in Appendix D. The lazy update mechanism used in our algorithm does not degrade the regret bound beyond a constant factor. This follows from the analysis of Abbasi-Yadkori et al. (2011), which shows that the elliptical potentials (related to the determinant of the design matrix) grow by at most a factor of $(1+C)$ compared to the standard design matrix of Algorithm 1. For the experimental design component, we adapt the proof of the standard elliptical lemma (Lattimore and Szepesvári, 2020, Lemma 19.4) to bring out the information gain, measured as $\log\det W_{t}$. This information gain is then related to $\log\det V_{t}$ thanks to standard results from submodular optimization. Specifically, the greedy strategy in Algorithm 3 achieves a $(1-1/e)$ approximation of the optimal information gain. Together, these results allow us to derive a regret bound for this improved algorithm, which matches the original up to constant factors.
5 Experiments
We validate our theoretical results on regret minimization (Theorems 3.4 and 4.1) on the Isaac-Cartpole-v0 environment from Nvidia Isaac Lab (Mittal et al., 2023). In this task the goal is to balance a pole on a cart by applying left or right forces, preventing the pole from falling. We compare RPO (for $\alpha=1$) with its lazy variants (LRPO without optimal design and LRPO-OD with optimal design). To simulate human preferences, we generate synthetic preferences using the built-in task-specific reward function. Furthermore, we adopt PPO (Schulman et al., 2017) as our RL oracle using 30 PPO steps per iteration of Algorithm 1 and 2. Moreover, to reduce variance in the reward estimate, we sample five trajectories from $\pi_{t}$ and $\pi_{t}^{\prime}$ in each RLHF round.
Figure 1: Comparison of RLHF algorithms in terms of (a) the ground truth reward $V_{\theta^{*}}^{\pi_{t}}$ (estimated from samples) and (b) number of preference queries performed. In particular, for the choice $\alpha=1$, we compare RPO (green, Algorithm 1) with it’s lazy versions LRPO and LRPO-OD (orange & blue). Here, LRPO and LRPO-OD refer to Algorithm 2 without and with optimal design subroutine. The solid lines indicate the median and the shaded areas the 0.2 and 0.8 quantiles, across 10 independent runs. The dashed blue line indicates the mean reward achieved by PPO with the ground truth parameter $\theta^{*}$.
Figure 1 shows that all three algorithms achieve performance competitive with RL using ground truth rewards. However, LRPO-OD reaches the highest performance while requiring significantly fewer preference queries. This highlights that, despite theoretical worst-case lower bounds, the number of preference queries can be considerably reduced in practice by selecting informative queries with optimal design. Additional experimental results and further implementation and evaluation details are provided in Appendix F. Furthermore, our code to run these experiments on Isaac Lab’s manager-based environments is included in the supplementary materials and will be publicly released.
6 Conclusion
We introduced a simple meta-algorithm for reinforcement learning from human feedback (RLHF) that leverages randomized exploration to achieve both regret and PAC-style guarantees. Our framework applies to general MDPs and utilizes an RL oracle as a subroutine. Notably, to our knowledge, our PAC-RL guarantee is the first for a randomized algorithm in both RLHF and RL settings. Building on this foundation, we developed an improved variant with better scalability and query efficiency, for which we also established regret guarantees. This variant employs lazy updates to enable parallelization of preference oracle calls, significantly reducing the demand on human annotators. Additionally, we apply greedy optimal design to select informative comparisons. Empirically, our approach is competitive with reward-based RL while requiring significantly fewer preference queries. Overall, our contributions advance the state of RLHF by combining strong theoretical guarantees with practical algorithm design, improving efficiency and broadening applicability to real-world scenarios.
Our work opens several directions for future research. First, we provide separate algorithms for regret minimization and PAC guarantees; it remains unclear whether this separation is necessary, and developing a unified algorithm that simultaneously achieves both objectives is an open challenge. Second, we adopt the Bradley-Terry model for preference generation, which may not fully capture the complexity of real-world human feedback. Extending the framework to richer preference models is an important direction. Third, our approach relies on an RL oracle at every RLHF step, which may be computationally demanding. While using a reward-free algorithm as an RL oracle is theoretically efficient, practical RL implementations are typically based on policy optimization, which is not a reward-free algorithm, and often entails high sample complexity. Thus, it remains open whether we could require fewer RL oracle calls or whether reward-free oracles can be successfully implemented. Finally, our experimental evaluation is limited to simple robotic control tasks with synthetic feedback. Assessing performance on more complex tasks and with real human or LLM-generated feedback would offer a stronger test of the method’s practical applicability.
Appendix
[…]
Appendix A Technical results
This section presents the technical results necessary for our theorems’ proofs.
A.1 Confidence set
The first result is a confidence set for the maximum likelihood estimation. It is an elliptical relaxation of the confidence set provided in Theorem 3.1 of Lee et al. [2024].
Lemma A.1 (Confidence set for the reward estimation).
Define for a given $\lambda>0$ the following set:
$\mathcal{E}_{t}(\delta):=\left\{{\theta:\left\lVert\theta-\hat{\theta}_{t}\right\rVert_{V_{t}}^{2}\leq\beta_{t}(\delta)^{2}:=\kappa\left[{\log\left({\frac{1}{\delta}}\right)+d\log\left({\max\left\{e,\frac{4eBLH_{\gamma}(t-1)}{d}\right\}}\right)}\right]+4\lambda B^{2}}\right\},$
with $V_{t}=\sum_{s=1}^{t-1}x_{t}x_{t}^{\top}+\lambda I$, $\kappa=\max_{\left\lVert\theta\right\rVert\leq B,x\in\mathcal{X}}\frac{1}{\dot{\sigma}(\langle\theta,x\rangle)}=\mathcal{O}(e^{2BLH_{\gamma}})$, $H_{\gamma}=(1-\gamma)^{-1}$, and $\dot{\sigma}$ is the first derivative of ${\sigma}$.
Then, it holds that:
$\Pr\left[{\exists t\geq 1:\theta^{*}\notin\mathcal{E}_{t}(\delta)}\right]\leq\delta,$
that is, $\mathcal{E}_{t}(\delta)$ is a high-probability confidence set for $\theta^{*}$ uniformly in time.
Proof.
By a first-order Taylor approximation with integral remainder, we have
$\mathcal{L}_{\mathcal{D}_{t}}(\theta^{*})=\mathcal{L}_{\mathcal{D}_{t}}(\hat{\theta}_{t})+\langle\nabla\mathcal{L}_{\mathcal{D}_{t}}(\hat{\theta}_{t}),\theta^{*}-\hat{\theta}_{t}\rangle+(\theta^{*}-\hat{\theta}_{t})^{\top}G_{t}(\hat{\theta}_{t},\theta^{*})(\theta^{*}-\hat{\theta}_{t}),$
where $\mathcal{L}_{\mathcal{D}_{t}}$ was defined in Equation (2) and
$G_{t}(\hat{\theta}_{t},\theta^{*})$ $=\int_{0}^{1}(1-\tau)\left({\sum_{s=1}^{t-1}\dot{\sigma}(\langle\hat{\theta}_{t}+\tau(\theta^{*}-\hat{\theta}_{t}),x_{s}\rangle)x_{s}x_{s}^{\top}}\right)\mathop{}\!\mathrm{d}\tau$
$=\sum_{s=1}^{t-1}\left[{\int_{0}^{1}(1-\tau)\dot{\sigma}(\langle\hat{\theta}_{t}+\tau(\theta^{*}-\hat{\theta}_{t}),x_{s}\rangle)\mathop{}\!\mathrm{d}\tau}\right]x_{s}x_{s}^{\top}$
$\succeq\kappa^{-1}\sum_{s=1}^{t-1}x_{s}x_{s}^{\top}.$ (3)
Rearranging terms gives
$\mathcal{L}_{\mathcal{D}_{t}}(\theta^{*})-\mathcal{L}_{\mathcal{D}_{t}}(\hat{\theta}_{t})$ $\stackrel{{\scriptstyle(i)}}{{\geq}}(\theta^{*}-\hat{\theta}_{t})^{\top}G_{t}(\hat{\theta}_{t},\theta^{*})(\theta^{*}-\hat{\theta}_{t})$
$\stackrel{{\scriptstyle(ii)}}{{\geq}}(\theta^{*}-\hat{\theta}_{t})^{\top}\left({\kappa^{-1}\sum_{s=1}^{t-1}x_{s}x_{s}^{\top}}\right)(\theta^{*}-\hat{\theta}_{t})$
$=\kappa^{-1}\left\lVert\theta^{*}-\hat{\theta}_{t}\right\rVert_{V_{t}}^{2}-\kappa^{-1}\lambda\left\lVert\theta^{*}-\hat{\theta}_{t}\right\rVert^{2},$
[…]
$\left\lVert\theta^{*}-\hat{\theta}_{t}\right\rVert_{V_{t}}^{2}\leq\kappa\left[{\log\left({\frac{1}{\delta}}\right)+d\log\left({\max\left\{{e,\frac{4eBLH_{\gamma}(t-1)}{d}}\right\}}\right)}\right]+4\lambda B^{2}.$
A.2 Optimism with constant probability
We first recall the following standard concentration and anti-concentration property of the Gaussian distribution.
Lemma A.2 (Appendix A of [Abeille and Lazaric, 2017]).
Let $z\sim\mathcal{N}(0,I)$ be a $d$-dimensional Gaussian random vector. Then, we have:
  1. 1.
    Anti-concentration: For any $u\in\mathcal{B}^{d}(1)$, we have $\Pr\left[{\langle u,z\rangle\geq 1}\right]\geq\frac{1}{4\sqrt{e\pi}}$.
  2. 2.
    Concentration: $\Pr\left[{\left\lVert z\right\rVert\leq\sqrt{2d\log(2d/\delta)}}\right]\geq 1-\delta$.
The anti-concentration property yields the following key result, which is required to prove a constant probability of optimism and subsequently control pessimism terms in the regret. Although it has been proven by Abeille and Lazaric [2017] in their linear Thompson sampling analysis, we provide a concise proof based on convex analysis for completeness.
Lemma A.3.
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a continuous and convex function, and consider the ellipsoid :=$\mathcal{E}:=\left\{{x\in\mathbb{R}^{d}:\left\lVert x-x_{0}\right\rVert_{A}\leq b}\right\}$ for a positive definite matrix $A$ and $b>0$. If $\tilde{x}\sim\mathcal{N}(x_{0},b^{2}A^{-1})$, then $\Pr\left[{f(\tilde{x})\geq\max_{x\in\mathcal{E}}f(x)}\right]\geq 1/\left({4\sqrt{e\pi}}\right)$.
Proof.
Note that by definition of $\tilde{x}$ we have $\tilde{x}\stackrel{{\scriptstyle d}}{{=}}x_{0}+bA^{-1/2}\tilde{z}$, where $\tilde{z}\sim\mathcal{N}(0,I)$. Hence, considering $g(z):=f(x_{0}+bA^{-1/2}z)$, we have
$p:=\Pr\left[{f(\tilde{x})\geq\max_{x\in\mathcal{E}}f(x)}\right]=\Pr\left[{g(\tilde{z})\geq\max_{z\in\mathcal{B}^{d}(1)}g(z)}\right],$
where we used that $x_{0}+bA^{-1/2}z\in\mathcal{E}$ if and only if $z\in\mathcal{B}^{d}(1)$. Since $g$ is a continuous convex function, we can choose $\bar{z}\in\operatorname*{arg\,max}_{z\in\mathcal{B}^{d}(1)}g(z)$ such that $\left\lVert\bar{z}\right\rVert=1$. By optimality of $\bar{z}$ it holds that [Rockafellar, 1997, Theorem 32.4]
$\partial f(\bar{z})\subseteq N_{\mathcal{B}^{d}(1)}(\bar{z}),$
where $N_{\mathcal{B}^{d}(1)}(\bar{z})=\left\{{h\in\mathbb{R}^{d}:h=\lambda\bar{z},\lambda\geq 0}\right\}$ denotes the normal cone to $\mathcal{B}^{d}(1)$ at $\bar{z}$. Hence, by convexity of $g$ we have
$g(\tilde{z})\geq g(\bar{z})+\langle\lambda\bar{z},\tilde{z}-\bar{z}\rangle=g(\bar{z})+\lambda(\langle\bar{z},\tilde{z}\rangle-1),\quad\text{ for some }\lambda\geq 0.$
Therefore, $\langle\bar{z},\tilde{z}\rangle\geq 1$ implies that $g(\tilde{z})\geq g(\bar{z})$, which yields the lower bound
$p=\Pr\left[{g(\tilde{z})\geq g(\bar{z})}\right]\geq\Pr\left[{\langle\tilde{z},\bar{z}\rangle\geq 1}\right].$
In light of Lemma A.2, this implies that $p\geq 1/(4\sqrt{e\pi})$. ∎
A.3 Elliptical potential bounds
The following lemma provides an upper bound on the sum of norms of sequentially observed vectors in the norm induced by their design matrix. These norms of vectors are commonly called elliptical potentials.
Lemma A.4 (Lemma 19.4 of Lattimore and Szepesvári [2020]).
Let $\{x_{t}\}_{t\geq 0}\in\mathbb{R}^{d}$ and for all $t\geq 0$, $\left\lVert x_{t}\right\rVert\leq L$, let $V_{t}=\lambda I+\sum_{s=0}^{t-1}x_{s}x_{s}^{\top}$ for some $\lambda>0$. Then,
$\sum_{t=1}^{T}\min\{1,\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}\}\leq\>$ $2d\log\left(1+\frac{TL^{2}}{d\lambda}\right).$
In the analysis of our algorithms, a central challenge is to control the norms of policy feature differences $\phi(\pi_{t})-\phi(\pi_{t})^{\prime}$. However, the learner only observes the trajectory features $\phi(\tau_{t})$ and $\phi(\tau_{t}^{\prime})$, which are random realizations of the policy features. To overcome this, we build on Lemma A.4 and introduce new tools to bound the sum of norms of policy feature differences.
Lemma A.5 (Elliptical lemma).
1) Let $\{x_{t}\}_{t\geq 0}\in\mathbb{R}^{d}$ and for all $t\geq 0$, $\left\lVert x_{t}\right\rVert\leq L$, let $V_{t}=\lambda I+\sum_{s=0}^{t-1}x_{s}x_{s}^{\top}$ for some $\lambda>0$. Then,
$\sum_{t=1}^{T}\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}\leq\>$ $2d\log\left(1+\frac{TL^{2}}{d\lambda}\right)+\frac{3dL^{2}}{\log(2)\lambda}\log\left(1+\frac{L^{2}}{\log(2)\lambda}\right).$
2) Let $\{x_{t}\}_{t\geq 0}\in\mathbb{R}^{d}$ be a sequence of random vectors adapted to a filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$. Assume that for all $t\geq 0$, $\left\lVert x_{t}\right\rVert\leq L$ almost surely. Then, for all $\delta>0$, it holds with probability at least $1-\delta$ that:
$\forall T\in\mathbb{N},\quad\sum_{t=1}^{T}\mathbb{E}[\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t}]\leq$ $2\sqrt{T\left({2d\log\left(1\!+\!\frac{TL^{2}}{d\lambda}\right)\!\!+\!\frac{3dL^{2}}{\log(2)\lambda}\log\left(1\!+\!\frac{L^{2}}{\log(2)\lambda}\right)}\right)}$
$+\frac{8L}{\sqrt{\lambda}}\log(1/\delta),$
where $V_{t}=\lambda I+\sum_{s=0}^{t-1}x_{s}x_{s}^{\top}$ for some $\lambda>0$.
The first statement is a small improvement over Lemma A.4 because it involves $\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}$ instead of $\min\{1,\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}\}$ and maintains a similar upper bound. In the second statement, $\{x_{t}\}_{t\geq 0}$ represent trajectory features and their expected values are policy features. Hence, the second statement of the lemma above allows us to control the elliptical potentials of the policy features while only observing trajectory features.
Proof.
First statement: The proof of this result is based on the observation in [Lattimore and Szepesvári, 2020, Exercise 19.3]. Namely, the number of times the term $\left\lVert x_{t}\right\rVert_{{V}_{t}^{-1}}^{2}$ can be larger than one is at most $\frac{3d}{\log(2)}\log(1+\frac{L^{2}}{\lambda\log(2)})$.
Let’s define the rounds $\mathcal{T}_{T}=\{t\leq T,\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}\leq 1\}$, we have:
$\sum_{t=1}^{T}\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}$ $=\sum_{t\in\mathcal{T}_{T}}\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}+\sum_{t\notin\mathcal{T}_{T}}\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}$
$\leq\sum_{t\in\mathcal{T}_{T}}\min\{1,\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}\}+\frac{3dL^{2}}{\lambda\log(2)}\log(1+\frac{L^{2}}{\lambda\log(2)}),$
where the first term of the last inequality follows by definition of $\mathcal{T}_{T}$. The second term follows because the number of times $1\leq t\leq T$ not in $\mathcal{T}_{T}$ is at most $\frac{3d}{\log(2)}\log(1+\frac{L^{2}}{\lambda\log(2)})$ as previously discussed, and because $\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}\leq L^{2}/\lambda$. Then, the proof is concluded by bounding the first sum on the right-hand side using Lemma A.4.
Second statement: The proof proceeds by using Lemma A.9. We have for any $\delta>0$ that with probability $1-\delta$:
$\sum_{t=1}^{T}\mathbb{E}[\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t}]$ $\leq 2\sum_{t=1}^{T}\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}+\frac{8L}{\sqrt{\lambda}}\log(1/\delta)$
$\leq 2\sqrt{T\sum_{t=1}^{T}\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}}+\frac{8L}{\sqrt{\lambda}}\log(1/\delta)$
$\leq 2\sqrt{T\left({2d\log\left(1\!+\!\frac{TL^{2}}{d\lambda}\right)\!\!+\!\frac{3dL^{2}}{\log(2)\lambda}\log\left(1\!+\!\frac{L^{2}}{\log(2)\lambda}\right)}\right)}+\frac{8L}{\sqrt{\lambda}}\log(1/\delta),$
where the first inequality uses Lemma A.9, the second uses the Cauchy-Schwarz inequality, and the last follows from the first result of the lemma. ∎
We now present a variant of the elliptical potentials lemma above, adapted for the case where the design matrix is updated with optimal design and lazy updates; see Algorithm 2 for more details.
Lemma A.6 (Lazy elliptical lemma).
Let $\{x_{t}\}_{t\geq 0}\in\mathbb{R}^{d}$ be a sequence of random vectors adapted to a filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$. Assume that for all $t\geq 0$, $\left\lVert x_{t}\right\rVert\leq L$. Then, we have the following results:
1) For all $\lambda>0$ and $C>0$, it holds that
$\sum_{t=1}^{T}\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}\leq\>$ $\frac{2(1+C)d}{1-1/e}\log\left(1+\frac{TL^{2}}{d\lambda}\right)+\frac{3dL^{2}}{\log(2)\lambda}\log\left(1+\frac{L^{2}}{\log(2)\lambda}\right).$
where $V_{t}$ is the matrix in Algorithm 2. Recall that $V_{t}$ is defined with the following lazy update procedure:
  • Initialize $V_{0}=\lambda I$, $W_{0}=\lambda I$, $\mathcal{D}=\emptyset$,
  • At a given round $t\in\mathbb{N}$:
    • Update $W_{t}$:
        If $\mathcal{D}=\emptyset$ update $W_{t}=V_{t}+x_{t-1}x_{t-1}^{\top}$,
        else $W_{t}=W_{t-1}+x_{t-1}x_{t-1}^{\top}$, $V_{t}=V_{t-1}$.
    • Update set $\mathcal{D}=\mathcal{D}\cup\{x_{t-1}\}$
    • if $\det W_{t}>(1+C)\det V_{t}$:
        update $V_{t}=V_{t-1}+\sum_{x\in\operatorname{D-OptDes}\left({\mathcal{D},\det(W_{t})}\right)}xx^{\top}$
        reset $\mathcal{D}=\emptyset$.
the operator $\operatorname{D-OptDes}$ is the procedure described in Algorithm 3.
2) For all $\delta>0$, it holds with probability at least $1-\delta$ that:
$\forall T\in\mathbb{N},\quad\sum_{t=1}^{T}\mathbb{E}[\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t}]\leq$ $\sqrt{\frac{8(1+C)dT}{(1-1/e)^{2}}\log\left(1+\frac{TL^{2}}{d{\lambda}}\right)\!+\!\frac{12dL^{2}T}{\log(2){\lambda}}\log\left(1\!+\!\frac{L^{2}}{\log(2){\lambda}}\right)}$
$+\frac{8L}{\sqrt{\lambda}}\log(1/\delta).$
Compared to Lemma A.5, we lose a factor of $\frac{1+C}{1-1/2}$ in the first statement and of $\frac{1+C}{(1-1/e)^{2}}$ in the second statement. The factor $(1+C)$ arises because we only update the matrix $V_{t}$ once $\det W_{t}$ exceeds $\det V_{t}$ by a constant $(1+C)$. The $1/(1-1/e)$ factors arise from the use of optimal design.
Proof.
First statement: We proceed in two steps: First, we relate $V_{t}$ to $W_{t}$, similar to the standard proofs for lazy updates (see proof of Theorem 4 of Abbasi-Yadkori et al. [2011] for example); The second step consists of using arguments of approximate optimality of greedy optimal design from [Nemhauser et al., 1978].
Step 1: Here, we relate the matrix norm in $V_{t}^{-1}$ to the matrix norm in $W_{t}$.
We have for all $t\in\mathbb{N}$ that $W_{t}\geq V_{t}$, which implies that $V_{t}^{-1}\geq W_{t}^{-1}$. Then, using Lemma A.8 we have:
$\forall x\in\mathbb{R}^{d},\quad\left\lVert x\right\rVert_{V_{t}^{-1}}^{2}\leq\left\lVert x\right\rVert_{W_{t}^{-1}}^{2}\frac{\det(V_{t}^{-1})}{\det(W_{t}^{-1})}.$ (4)
By definition of $V_{t}$, we know that for all $t\geq 1$, $\det(W_{t})\leq(1+C)\det(V_{t})$, which implies that $\det(V_{t}^{-1})\leq(1+C)\det(W_{t}^{-1})$. Then, from (4) we obtain that:
$\forall x\in\mathbb{R}^{d},\quad\left\lVert x\right\rVert_{V_{t}^{-1}}^{2}\leq(1+C)\left\lVert x\right\rVert_{W_{t}^{-1}}^{2}.$
Now, for rounds $t_{\operatorname{stop}}\in\mathbb{N}$ such that $\det V_{t_{\operatorname{stop}}}>(1+C)\det V_{t_{\operatorname{stop}}-1}$, and all rounds $t>t_{\operatorname{stop}}$ such that $\det W_{t}\leq(1+C)\det V_{t_{\operatorname{stop}}}$ we have:
$\sum_{q=t_{\operatorname{stop}}}^{t}\left\lVert x_{q}\right\rVert_{V_{t}^{-1}}^{2}=\sum_{q=t_{\operatorname{stop}}}^{t}\left\lVert x_{q}\right\rVert_{V_{t_{\operatorname{stop}}}^{-1}}^{2}\leq(1+C)\sum_{q=t_{\operatorname{stop}}}^{t}\left\lVert x_{q}\right\rVert_{W_{q}^{-1}}^{2}.$
In addition, we have:
$\sum_{q=t_{\operatorname{stop}}}^{t-1}\min\{1,\left\lVert x_{q}\right\rVert_{V_{t}^{-1}}^{2}\}\leq 2(1+C)\sum_{q=t_{\operatorname{stop}}}^{t-1}\log(1+\left\lVert x_{q}\right\rVert_{W_{q}^{-1}}^{2})=2(1+C)\log(\det(W_{t})/\det(W_{t_{\operatorname{stop}}})),$
where the last equality follows because $\frac{\det(W_{q+1})}{\det(W_{q})}=1+\left\lVert x_{q}\right\rVert_{W_{q}^{-1}}^{2}$ (see the proof of [Abbasi-Yadkori et al., 2011, Lemma 11] for example).
Step 2: We now relate the terms $(\log\det(W_{t}))_{t\leq t_{\operatorname{stop}}^{\prime}}$ to the term $\log\det(V_{t_{\operatorname{stop}}^{\prime}})$ where $t_{\operatorname{stop}}^{\prime}$ is the first time greater than $t_{\operatorname{stop}}$ such that $\det(V_{t_{\operatorname{stop}}^{\prime}})>(1+C)\det(V_{t_{\operatorname{stop}}^{\prime}-1})$.
Based on Algorithm 3, there are two possibilities:
  1. 1.
    either the new design matrix $V_{t_{\operatorname{stop}}^{\prime}}$ is such that $\det V_{t_{\operatorname{stop}}^{\prime}}>\det W_{t_{\operatorname{stop}}^{\prime}}$,
  2. 2.
    or the procedure $\operatorname{D-OptDes}\left({\mathcal{D},\det(W_{t_{\operatorname{stop}}^{\prime}})}\right)$ terminates after adding $|\mathcal{D}|$ elements to the matrix $V_{t_{\operatorname{stop}}}$.
In the first case, we have by definition that $\det V_{t_{\operatorname{stop}}^{\prime}}\geq\det W_{t_{\operatorname{stop}}^{\prime}}$. We now show that a similar result also holds in the second case.
Consider the set $\mathcal{D}$ at time $t_{\operatorname{stop}}^{\prime}$. The set function $S\subset\mathcal{D}\to\log\det\left({V_{t_{\operatorname{stop}}}+\sum_{x\in S}xx^{T}}\right)$ is submodular [Krause et al., 2008]. Therefore, it follows from [Nemhauser et al., 1978] that the greedy algorithm maximizing the above function is $(1-1/e)$ optimal. Namely, consider the set $\mathcal{D}_{\text{opt}}$ defined by greedily adding the features $x\in\mathcal{D}$ until $|\mathcal{D}_{\text{opt}}|=|\mathcal{D}|$. Then, it holds that:
$\log\det\left({V_{t_{\operatorname{stop}}}+\sum_{x\in\mathcal{D}_{\text{opt}}}xx^{T}}\right)\geq(1-1/e)\max_{S\subset\mathcal{D},|S|\leq|\mathcal{D}|}\log\det\left({V_{t_{\operatorname{stop}}}+\sum_{x\in S}xx^{T}}\right).$
From the above, we deduce that for all $t\leq t_{\operatorname{stop}}^{\prime}$:
$\log\det(W_{t})=\log\det(V_{t_{\operatorname{stop}}}+\sum_{s=t_{\operatorname{stop}}}^{t}x_{s}x_{s}^{T})$ $\leq\max_{S\subset\mathcal{D},|S|\leq|\mathcal{D}|}\log\det\left({V_{t_{\operatorname{stop}}}+\sum_{x\in S}xx^{T}}\right)$
$\leq\frac{1}{1-1/e}\log\det\left({V_{t_{\operatorname{stop}}}+\sum_{x\in\mathcal{D}_{\text{opt}}}xx^{T}}\right).$
We conclude that in both cases of termination of Algorithm 3, it holds for all $t\leq t_{\operatorname{stop}}^{\prime}$ that: $\log\det(W_{t})\leq\frac{1}{1-1/e}\log\det\left({V_{t_{\operatorname{stop}}^{\prime}}}\right)$.
Combination of steps 1 and 2: From step 1, we know that:
$\sum_{q=1}^{t-1}\min\{1,\left\lVert x_{q}\right\rVert_{V_{t}^{-1}}^{2}\}\leq 2(1+C)\sum_{q=1}^{t-1}\log(1+\left\lVert x_{q}\right\rVert_{W_{q}^{-1}}^{2})=2(1+C)\log(\det(W_{t})/\det(W_{0})),$
In addition, exercise 19.3 of [Lattimore and Szepesvári, 2020] proves that the number of times where $\left\lVert x_{q}\right\rVert_{W_{t}^{-1}}$ is greater than one is at most $\frac{3d}{\log(2)}\log\left({1+\frac{L^{2}}{\lambda\log(2)}}\right)$. Therefore, we deduce that for all $T\in\mathbb{N}$:
$\sum_{q=0}^{T}\left\lVert x_{q}\right\rVert_{V_{t}^{-1}}^{2}$ $\leq\sum_{q=0}^{T}\min\{1,\left\lVert x_{q}\right\rVert_{V_{t}^{-1}}^{2}\}+\frac{3dL^{2}}{\lambda\log(2)}\log\left({1+\frac{L^{2}}{\lambda\log(2)}}\right)$
$\leq 2(1+C)\log(\det(W_{t})/\det(W_{0}))+\frac{3dL^{2}}{\lambda\log(2)}\log\left({1+\frac{L^{2}}{\lambda\log(2)}}\right).$
Finally, plugging the result of step 2 above yields that:
$\sum_{q=0}^{T}\left\lVert x_{q}\right\rVert_{V_{t}^{-1}}^{2}\leq\frac{2(1+C)}{1-1/e}\log(\det(V_{T})/\det(V_{0}))+\frac{3dL^{2}}{\lambda\log(2)}\log\left({1+\frac{L^{2}}{\lambda\log(2)}}\right)$
$\leq\frac{2(1+C)d}{1-1/e}\log\left({1+\frac{TL^{2}}{\lambda}}\right)+\frac{3dL^{2}}{\lambda\log(2)}\log\left({1+\frac{L^{2}}{\lambda\log(2)}}\right).$
This concludes the proof of the first result.
Second statement: We know from inequality (4) that for all $t\in[\tau,\tau^{\prime}-1]$:
$\mathbb{E}[\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t}]\leq\mathbb{E}[\left\lVert x_{t}\right\rVert_{W_{t}^{-1}}|\mathcal{F}_{t}].$
Then, using Lemma A.9 we obtain:
$\sum_{t=\tau}^{\tau^{\prime}-1}\mathbb{E}[\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t}]\leq 2\sum_{t=\tau}^{\tau^{\prime}-1}\left\lVert x_{t}\right\rVert_{W_{t}^{-1}}+\frac{8L}{\sqrt{\lambda}}\log(1/\delta).$
Then, we can conclude the proof using the same arguments as for the first result. ∎
A.4 Miscellaneous
Proposition A.7.
The function $f(\theta)=\sup_{\pi\in\left({\Delta_{\mathcal{A}}}\right)^{\mathcal{S}}}\>\langle\theta,\phi(\pi)\rangle$ is convex and continuous over $\mathbb{R}^{d}$.
Proof.
The function $f$ is the support function of the set $\mathcal{Z}=\left\{{\phi(\pi):\pi:\mathcal{S}\to\Delta_{\mathcal{A}}}\right\}$. To prove the convexity, note that for any $\eta\in(0,1)$, we have
$f(\eta\theta+(1-\eta)\theta^{\prime})\leq\eta\sup_{z\in\mathcal{Z}}\langle\theta,z\rangle+(1-\eta)\sup_{z\in\mathcal{Z}}\langle\theta^{\prime},z\rangle\leq\eta f(\theta)+(1-\eta)f(\theta^{\prime}).$
Furthermore, any convex function is continuous over the relative interior of its effective domain $\operatorname{\operatorname{dom}}f=\left\{{x:f(x)<\infty}\right\}$ (see e.g. [Rockafellar, 1997, Theorem 10.1]). Since $|f(\theta)|\leq\left\lVert\theta\right\rVert LH_{\gamma}$, this implies that $f$ is continuous over $\mathbb{R}^{d}$. ∎
Lemma A.8 (Lemma 12 of [Abbasi-Yadkori et al., 2011]).
Let $A,B$, and $C$ be positive semi-definite matrices such that $A=B+C$. Then, we have that:
$\sup_{x\neq 0}\frac{x^{T}Ax}{x^{T}Bx}\leq\frac{\det(A)}{\det(B)}.$
This next lemma is one of the versions of Freedman’s inequality [Freedman, 1975].
Lemma A.9 (Lemma 2 of Zhu and Nowak [2022]).
Let $\left(z_{t}\right)_{t\leq T}$ be a real-valued sequence of random variables adapted to a filtration $\mathcal{F}_{t}$. If $\left|z_{t}\right|\leq B$ almost surely, then with probability at least $1-\delta$,
$\sum_{t=1}^{T}z_{t}\leq\frac{3}{2}\sum_{t=1}^{T}\mathbb{E}_{t}\left[z_{t}\right]+4B\log\left(2\delta^{-1}\right)$
and
$\sum_{t=1}^{T}\mathbb{E}_{t}\left[z_{t}\right]\leq 2\sum_{t=1}^{T}z_{t}+8B\log\left(2\delta^{-1}\right).$
Appendix B Proof of Theorem 3.4
Proof.
The proof proceeds in four steps. We first define a set of good events of concentration of parameters and sums of trajectory features, we show that they are satisfied with high probability. Second, we decompose the regret into two terms: a pessimism term and an estimation error term. We then show that the pessimism term is controlled by establishing a constant probability of optimism, and the estimation error term is controlled by the estimation error of the maximum likelihood estimator.
Throughout the proof, we work with the two filtrations $\mathcal{F}_{t}=\sigma(x_{1},y_{1},\ldots,x_{t-1},y_{t-1})$ and $\mathcal{F}_{t}^{\theta}=\sigma(\mathcal{F}_{t},\tilde{\theta}_{t})$. In particular, both $\hat{\theta}_{t}$ and $V_{t}$ are $\mathcal{F}_{t}$ measurable. Therefore, $\tilde{\theta}_{t}$ follows a Gaussian distribution given $\mathcal{F}_{t}$, i.e. $\tilde{\theta}_{t}\mid\mathcal{F}_{t}\sim\mathcal{N}(\hat{\theta}_{t},\beta_{t}^{2}V_{t}^{-1})$, while it is fully determined by $\mathcal{F}_{t}^{\theta}$.
Step 1 (good events): Recall that in Algorithm 1 we set $\varepsilon=1/\sqrt{T}$ and $\delta^{\prime}=\delta/5$. We define the following high probability events.
  1. 1.
    Let $\delta^{\prime\prime}=\delta^{\prime}/T$ and $c(\delta^{\prime\prime}):=\sqrt{2d\log(2d/\delta^{\prime\prime})}=\sqrt{2d\log(10dT/\delta)}$. Consider the inflated ellipsoid
    $\mathcal{E}_{t}^{\text{TS}}:=\left\{{\theta\in\mathbb{R}^{d}:\left\lVert\theta-\hat{\theta}_{t}\right\rVert_{V_{t}}\leq\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})}\right\}=c(\delta^{\prime\prime})\mathcal{E}_{t}(\delta^{\prime}).$
    We define the events
    $\widehat{E}_{t}:=\left\{{\theta^{*}\in\mathcal{E}_{t}}\right\},\quad\widetilde{E}_{t}:=\left\{{(1-\alpha)\hat{\theta}_{t}+\tilde{\theta}_{t}\in\mathcal{E}_{t}^{\text{TS}}}\right\},\quad\text{ and }E_{t}:=\widehat{E}_{t}\cap\widetilde{E}_{t}$
    and let $G_{1}:=\bigcap_{t=1}^{T}E_{t}$. This event implies that $\theta^{*}$ lies in the confidence set uniformly over times $t\leq T$, and the sampled parameter is close to $\alpha\hat{\theta}_{t}$.
  2. 2.
    Let $G_{2}$ denote the event that for
    $C_{T}:=2\sqrt{T\left({2d\log\left(1\!+\!\frac{4TL^{2}H_{\gamma}^{2}}{d\lambda}\right)\!\!+\!\frac{12dL^{2}H_{\gamma}^{2}}{\log(2)\lambda}\log\left(1\!+\!\frac{4L^{2}H_{\gamma}^{2}}{\log(2)\lambda}\right)}\right)}+\frac{16LH_{\gamma}}{\sqrt{\lambda}}\log\left({\frac{2}{\delta^{\prime}}}\right),$
    it holds that
    $\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t}}\right]\leq C_{T},$
    and that
    $\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t}^{\theta}}\right]\leq C_{T}.$
  3. 3.
    Let $G_{3}=\left\{{\forall t\in[1,T]:V_{\tilde{\theta}_{t}}^{*}-V_{\tilde{\theta}_{t}}^{\pi_{t}}\leq\varepsilon}\right\}$, where $\pi_{t}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\tilde{\theta}_{t},\varepsilon,\delta^{\prime}/T}\right)$.
  4. 4.
    Let $G_{4}$ denote the event that $V_{\bar{\theta}_{T}}^{*}-V_{\bar{\theta}_{T}}^{\hat{\pi}}\leq\varepsilon$ where $\hat{\pi}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\bar{\theta}_{T},\varepsilon,\delta^{\prime}}\right)$.
As shown below, with probability at least $1-\delta$, all the above good events happen at the same time.
Proposition B.1.
Let $G:=\bigcap_{i=1}^{4}G_{i}$. It holds that $\Pr\left[{G}\right]\geq 1-\delta$.
Proof.
For the event $G_{1}$, Lemma 3.1 and A.2 imply that $\Pr\left[{\bigcup_{t=1}^{T}\widehat{E}_{t}^{\complement}}\right]\leq\delta^{\prime}$ and $\Pr[\widetilde{E}_{t}^{\complement}]\leq\delta^{\prime}/T$ for any $t\in[1,T]$. Hence, a union bound yields $\Pr\left[{G_{1}^{\complement}}\right]\leq 2\delta^{\prime}$. Furthermore, by Lemma A.5 2), we have $\Pr\left[{G_{2}^{\complement}}\right]\leq\delta^{\prime}$. Moreover, by union bound and the definition of $\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}$, we have $\Pr\left[{G_{3}^{\complement}}\right]\leq\delta^{\prime}$ and $\Pr\left[{G_{4}^{\complement}}\right]\leq\delta^{\prime}$. Hence, we have $\Pr\left[{G}\right]\geq 1-\sum_{i=1}^{4}\Pr\left[{G_{i}^{\complement}}\right]\geq 1-\delta$. ∎
Remark B.2.
Note that throughout this proof, we consider $\alpha=1$. However, we defined the good events for general $\alpha\in\left\{{0,1}\right\}$ so we can reuse Proposition B.1 for the proof of Theorem 3.5, which requires $\alpha=0$ instead. In particular, the event $G_{4}$ is relevant only for the setting with $\alpha=0$.
Step 2 (regret decomposition): Recall that in algorithm 1, we choose $\pi_{t}^{\prime}=\pi_{t-1}$. Hence, we can upper bound the cumulative regret as follows
$R(T)$ $=\frac{1}{2}\sum_{t=1}^{T}\left({2V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}}-V_{\theta^{*}}^{\pi^{\prime}_{t}}}\right)$
$\leq\sum_{t=1}^{T}\left({V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}}}\right)+V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{0}^{\prime}}$
$\leq\sum_{t=1}^{T}\underbrace{\left({V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}}}\right)}_{r_{t}}+BLH_{\gamma}.$
Let $\Delta_{t}(\theta):=\max_{\pi}\langle\theta,\phi(\pi)-\phi(\pi^{\prime}_{t})\rangle$. On the good event $G$, we can decompose the instantaneous regret as follows
$r_{t}$ $:=V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}}$
$=\left({V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}^{\prime}}}\right)-\left({V_{\tilde{\theta}_{t}}^{\pi_{t}}-V_{\tilde{\theta}_{t}}^{\pi_{t}^{\prime}}}\right)+\left({V_{\tilde{\theta}_{t}}^{\pi_{t}}-V_{\tilde{\theta}_{t}}^{\pi_{t}^{\prime}}}\right)-\left({V_{\theta^{*}}^{\pi_{t}}-V_{\theta^{*}}^{\pi_{t}^{\prime}}}\right)$
$\leq\left({V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}^{\prime}}}\right)-\left({V_{\tilde{\theta}_{t}}^{*}-V_{\tilde{\theta}_{t}}^{\pi_{t}^{\prime}}}\right)+\varepsilon+\left({V_{\tilde{\theta}_{t}}^{\pi_{t}}-V_{\tilde{\theta}_{t}}^{\pi_{t}^{\prime}}}\right)-\left({V_{\theta^{*}}^{\pi_{t}}-V_{\theta^{*}}^{\pi_{t}^{\prime}}}\right)$
$=\underbrace{\Delta_{t}(\theta^{*})-\Delta_{t}(\tilde{\theta}_{t})}_{r_{t}^{\text{TS}}}+\underbrace{\langle\tilde{\theta}_{t}-\theta^{*},\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\rangle}_{r_{t}^{\text{MLE}}}+\varepsilon.$
Here, $r_{t}^{\text{TS}}$ is a pessimism term that is negative by construction for optimistic algorithms, and $r_{t}^{\text{MLE}}$ is related to the estimation error of the reward parameter.
Step 3 (bounding ${r_{t}^{\text{TS}}}$): This part of the analysis highlights the distinctiveness of randomized exploration. While optimistic algorithms ensure negativity of ${r_{t}^{\text{TS}}}$ through their intractable optimization procedures, randomized exploration controls it using probability arguments. Specifically, following the proof of Abeille and Lazaric [2017], we begin by bounding $r_{t}^{\text{TS}}$ on the good event via a conditional expectation given the optimism event. The bound on ${r_{t}^{\text{TS}}}$ then follows by a careful application of the anti-concentration property established in Lemma A.3.
Conditioned on $\widetilde{E}_{t}$, we can lower bound
$\Delta_{t}(\tilde{\theta}_{t})\geq\min_{\theta\in\mathcal{E}_{t}^{\text{TS}}}\Delta_{t}(\theta)=\max_{\pi}\langle\underline{\theta}_{t},\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle=:\underline{\Delta}_{t},$
for some $\underline{\theta}_{t}\in\mathcal{E}^{\text{TS}}_{t}$. Moreover, if $O_{t}:=\left\{{\Delta_{t}(\tilde{\theta}_{t})\geq\Delta_{t}(\theta^{*})}\right\}$ denotes the event of $\tilde{\theta}_{t}$ being optimistic at time $t$, we can upper bound $\Delta_{t}(\theta^{*})$ as follows
$\Delta_{t}(\theta^{*})\mathbbm{1}(E_{t})\leq\mathbb{E}\left[{\Delta_{t}(\tilde{\theta}_{t})\mathbbm{1}(E_{t})\mid\mathcal{F}_{t},O_{t}}\right].$
Putting this together, we get
$r_{t}^{\text{TS}}\mathbbm{1}(E_{t})$ $\leq\mathbb{E}\left[{\left({\Delta_{t}(\tilde{\theta}_{t})-\underline{\Delta}_{t}}\right)\mathbbm{1}(E_{t})\mid\mathcal{F}_{t},O_{t}}\right]$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\mathbb{E}\left[{\left({\langle\tilde{\theta}_{t},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle+\varepsilon-\max_{\pi}\langle\underline{\theta}_{t},\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle}\right)\mathbbm{1}(E_{t})\mid\mathcal{F}_{t},O_{t}}\right]$
=$\stackrel{{\scriptstyle(ii)}}{{=}}\mathbb{E}\left[{\left({\langle\tilde{\theta}_{t},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle-\langle\underline{\theta}_{t},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle}\right)\mathbbm{1}(E_{t})\mid\mathcal{F}_{t},O_{t}}\right]+\varepsilon$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t},\widehat{E}_{t},O_{t}}\right]\Pr\left[{\widehat{E}_{t}\mid\mathcal{F}_{t}}\right]+\varepsilon,$
where $(i)$ follows from $\varepsilon$-optimality of $\pi_{t}$, in $(ii)$ we used that $\max_{\pi}\langle\underline{\theta}_{t},\phi(\pi)-\phi(\pi_{t})\rangle\geq 0$, and $(iii)$ follows from the Cauchy-Schwarz inequality. By the law of total probability, we have that
$\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t},\widehat{E}_{t},O_{t}}\right]\leq\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t},\widehat{E}_{t}}\right]/\Pr\left[{O_{t}|\mathcal{F}_{t},\widehat{E}_{t}}\right].$
Next, as $\theta^{*}\in\mathcal{E}_{t}$ on $\widehat{E}_{t}$, we have
$\Pr\left[{O_{t}\mid\mathcal{F}_{t},\widehat{E}_{t}}\right]\geq\Pr\left[{\Delta_{t}(\tilde{\theta}_{t})\geq\max_{\theta\in\mathcal{E}_{t}}\Delta_{t}(\theta)\mid\mathcal{F}_{t}}\right].$
Since $\theta\mapsto\Delta_{t}(\theta)$ is the sum of a linear function and the function $\theta\mapsto\max_{\pi}\langle\theta,\phi(\pi)\rangle$ which is convex and continuous by Proposition A.7. Then, applying Lemma A.3 with $f(\theta)=\Delta_{t}(\theta)$, $\tilde{x}=\tilde{\theta}_{t}\mid\mathcal{F}_{t}$, and $\mathcal{E}=\mathcal{E}_{t}$, we have
$\Pr\left[{O_{t}\mid\mathcal{F}_{t},\widehat{E}_{t}}\right]\geq 1/\left({4\sqrt{e\pi}}\right)=:p.$
As a result, we can upper bound the instantaneous regret as
$r_{t}^{\text{TS}}\mathbbm{1}(E_{t})$ $\leq\frac{2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t}}\right]}{p}+\varepsilon$
$\leq\frac{2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t}}\right]}{p}+\varepsilon,$
where the second inequality follows from the convexity of norms. Applying Lemma A.5 2), we have on the good event $G$ that
$\sum_{t=1}^{T}r_{t}^{\text{TS}}$ $\leq\dfrac{2\beta_{T}(\delta^{\prime})c(\delta^{\prime\prime})}{p}\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t}}\right]+T\varepsilon$ (5)
$\leq\dfrac{2\beta_{T}(\delta^{\prime})c(\delta^{\prime\prime})}{p}C_{T}+T\varepsilon,$
for the constant
$C_{T}=2\sqrt{T\left({2d\log\left(1\!+\!\frac{4TL^{2}H_{\gamma}^{2}}{d\lambda}\right)\!\!+\!\frac{12dL^{2}H_{\gamma}^{2}}{\log(2)\lambda}\log\left(1\!+\!\frac{4L^{2}H_{\gamma}^{2}}{\log(2)\lambda}\right)}\right)}+\frac{16LH_{\gamma}}{\sqrt{\lambda}}\log\left({\frac{2}{\delta^{\prime}}}\right).$
Step 4 (bounding ${r_{t}^{\text{MLE}}}$): Conditioned on event $E_{t}$ we have
$r_{t}^{\text{MLE}}$ $=\langle\tilde{\theta}_{t}-\theta^{*},\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\rangle$
$=\langle\tilde{\theta}_{t}-\hat{\theta}_{t},\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\rangle+\langle\hat{\theta}_{t}-\theta^{*},\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\rangle$
$\leq\beta_{t}(\delta^{\prime})(1+c(\delta^{\prime\prime}))\left\lVert\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\right\rVert_{V_{t}^{-1}}$
$\leq\beta_{t}(\delta^{\prime})(1+c(\delta^{\prime\prime}))\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t}^{\theta}}\right],$
where the last inequality follows from the convexity of norms. From here, we can again apply the second result of Lemma A.5. We deduce that on the good event $G$, we have
$\sum_{t=1}^{T}r_{t}^{\text{MLE}}$ $\leq\beta_{T}(\delta^{\prime})(1+c(\delta^{\prime\prime}))\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t}^{\theta}}\right]$
$\leq\beta_{T}(\delta^{\prime})(1+c(\delta^{\prime\prime}))C_{T}.$
In summary, we can conclude that with probability at least $1-\delta$, the regret can be bounded as follows
$R(T)$ $\leq\sum_{t=1}^{T}\left({r_{t}^{\text{TS}}+r_{t}^{\text{MLE}}}\right)+T\varepsilon+BLH_{\gamma}$
$\leq\left({\dfrac{2\beta_{T}(\delta^{\prime})c(\delta^{\prime\prime})}{p}+\beta_{T}(\delta^{\prime})(1+c(\delta^{\prime\prime}))}\right)C_{T}+2T\varepsilon+BLH_{\gamma}$
$\leq\dfrac{3\beta_{T}(\delta^{\prime})c(\delta^{\prime\prime})}{p}C_{T}+2T\varepsilon+BLH_{\gamma}$
$=\dfrac{3\left[{\sqrt{\kappa\left[{\log\left({\frac{5}{\delta}}\right)+d\log\left({\max\left\{{e,\frac{4eBLH_{\gamma}(T-1)}{d}}\right\}}\right)}\right]}+2\sqrt{\lambda}B}\right]\sqrt{2d\log(10dT/\delta)}}{p}$
$\cdot\left[{2\sqrt{T\left({2d\log\left(1\!+\!\frac{4TL^{2}H_{\gamma}^{2}}{d\lambda}\right)\!\!+\!\frac{12dL^{2}H_{\gamma}^{2}}{\log(2)\lambda}\log\left(1\!+\!\frac{4L^{2}H_{\gamma}^{2}}{\log(2)\lambda}\right)}\right)}+\frac{16LH_{\gamma}}{\sqrt{\lambda}}\log\left({\frac{10}{\delta}}\right)}\right]$
$+2\sqrt{T}+BLH_{\gamma}$
$=\mathcal{O}\left({\sqrt{\kappa d^{3}T\log\left({\dfrac{dT}{\delta}}\right)^{3}}}\right).$
Appendix C Proof of Theorem 3.5
Proof.
Recall that in Algorithm 1, we set $\varepsilon=1/\sqrt{T}$, $\delta^{\prime}=\delta/5$, and $\hat{\pi}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}(\bar{\theta}_{T},\varepsilon,\delta^{\prime})$ for $\bar{\theta}_{T}=\frac{1}{T}\sum_{t=1}^{T}\hat{\theta}_{t}$. Moreover, as $\alpha=0$, we further have $\tilde{\theta}_{t}\mid\mathcal{F}_{t}\sim\mathcal{N}(0,\beta_{t}^{2}V_{t}^{-1})$, and choose the policies $\pi_{t}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}(\tilde{\theta}_{t})$ and $\pi_{t}^{\prime}=\pi_{t-1}$.
In the following, we will denote $\pi^{*}\in\operatorname*{arg\,max}_{\pi}\langle\theta^{*},\phi(\pi)\rangle$ for an arbitrary optimal policy corresponding to the ground truth reward parameter $\theta^{*}$. Under the good event $G$ of Proposition B.1, we can bound the suboptimality of $\hat{\pi}$ as follows
$\mathsf{SubOpt}(\hat{\pi})$ $=\langle\theta^{*},\phi(\pi^{*})-\phi(\hat{\pi})\rangle$
$\leq\langle\theta^{*}-\bar{\theta}_{T},\phi(\pi^{*})-\phi(\hat{\pi})\rangle+\varepsilon$
$=\varepsilon+\frac{1}{T}\sum_{t=1}^{T}\langle\theta^{*}-\hat{\theta}_{t},\phi(\pi^{*})-\phi(\hat{\pi})\rangle,$
where we used $\varepsilon$-optimality of $\hat{\pi}$ and the definition of $\bar{\theta}_{T}$. Now, recall that conditioned on $E_{t}$, we have $\theta^{*}\in\mathcal{E}_{t}$ and $\hat{\theta}_{t}+\tilde{\theta}_{t}\in{\mathcal{E}}_{t}^{\text{TS}}$. Define ¯$\bar{\mathcal{E}}_{t}=\{\theta:\theta+\hat{\theta}\in\mathcal{E}_{t}\}$. Then, on the good event $G$ it holds that:
$\langle\theta^{*}-\hat{\theta}_{t},\phi(\pi^{*})-\phi(\hat{\pi})\rangle\mathbbm{1}\left({E_{t}}\right)$ $\leq\max_{\pi,\pi^{\prime}}\max_{\theta\in\mathcal{E}_{t}}\langle\theta-\hat{\theta}_{t},\phi(\pi)-\phi(\pi^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}}\right)$
$=\max_{\pi,\pi^{\prime}}\max_{\theta\in\bar{\mathcal{E}}_{t}}\langle\theta,\phi(\pi)-\phi(\pi^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}}\right)$
=$\stackrel{{\scriptstyle(i)}}{{=}}\beta_{t}\max_{\pi,\pi^{\prime}}\left\lVert\phi(\pi)-\phi(\pi^{\prime})\right\rVert_{V_{t}^{-1}}\mathbbm{1}\left({\widetilde{E}_{t}}\right)$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\beta_{t}\max_{\pi,\pi^{\prime}}\left({\left\lVert\phi(\pi)-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}+\left\lVert\phi(\pi_{t}^{\prime})-\phi(\pi^{\prime})\right\rVert_{V_{t}^{-1}}}\right)\mathbbm{1}\left({\widetilde{E}_{t}}\right)$
$=2\beta_{t}\max_{\pi}\left\lVert\phi(\pi)-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mathbbm{1}\left({\widetilde{E}_{t}}\right)$
=$\stackrel{{\scriptstyle(iii)}}{{=}}2\max_{\pi}\max_{\theta\in\bar{\mathcal{E}}_{t}}\langle\theta,\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}}\right)$
$=2\max_{\theta\in\bar{\mathcal{E}}_{t}}\max_{\pi}\langle\theta,\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}}\right).$
Here, $(i)$ and $(iii)$ follow because ¯$\bar{\mathcal{E}}_{t}$ is a centred ellipsoid and $(ii)$ from the triangle inequality. Next, we define the non-negative function $f_{t}(\theta):=\max_{\pi}\langle\theta,\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle$ and the event $O_{t}=\{f_{t}(\tilde{\theta}_{t})\geq\max_{\theta\in\bar{\mathcal{E}}_{t}}f_{t}(\theta)\}$. By Proposition A.7 $f_{t}$ is convex and continuous. Applying Lemma A.3 with $f=f_{t}$, $\tilde{x}=\tilde{\theta}_{t}\mid\mathcal{F}_{t}$, and =¯$\mathcal{E}=\bar{\mathcal{E}}_{t}$, it holds that $\Pr\left[{O_{t}\mid\mathcal{F}_{t}}\right]\geq p:=1/(4\sqrt{e\pi})$. Hence,
$\langle\theta^{*}-\hat{\theta}_{t},\phi(\pi^{*})-\phi(\hat{\pi})\rangle\mathbbm{1}\left({E_{t}}\right)$ $\leq 2\max_{\theta\in\bar{\mathcal{E}}_{t}}\max_{\pi}\langle\theta,\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}}\right)$
$\leq 2\mathbb{E}\left[{\max_{\pi}\langle\tilde{\theta}_{t},\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}}\right)\mid O_{t},\mathcal{F}_{t}}\right]$
$\leq 2\mathbb{E}\left[{\langle\tilde{\theta}_{t},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}}\right)\mid O_{t},\mathcal{F}_{t}}\right]+2\varepsilon$
$\leq 2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mid O_{t},\mathcal{F}_{t}}\right]+2\varepsilon$
$\leq\frac{2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})}{p}\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t}}\right]+2\varepsilon.$
Using Lemma A.5, we conclude that with probability $1-\delta$, we have that
$\mathsf{SubOpt}(\hat{\pi})\leq 3\varepsilon+\frac{2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})}{pT}\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t}}\right]$
$\leq$ $\frac{3}{\sqrt{T}}+\dfrac{2\left[{\sqrt{\kappa\left[{\log\left({\frac{5}{\delta}}\right)+d\log\left({\max\left\{{e,\frac{4eBLH_{\gamma}(T-1)}{d}}\right\}}\right)}\right]}+2\sqrt{\lambda}B}\right]\sqrt{2d\log(10dT/\delta)}}{pT}$
$\cdot\left[{2\sqrt{T\left({2d\log\left(1\!+\!\frac{4TL^{2}H_{\gamma}^{2}}{d\lambda}\right)\!\!+\!\frac{12dL^{2}H_{\gamma}^{2}}{\log(2)\lambda}\log\left(1\!+\!\frac{4L^{2}H_{\gamma}^{2}}{\log(2)\lambda}\right)}\right)}+\frac{16LH_{\gamma}}{\sqrt{\lambda}}\log\left({\frac{10}{\delta}}\right)}\right]$
$=$ $\mathcal{O}\left({\sqrt{\dfrac{\kappa d^{3}}{T}\log\left({\dfrac{dT}{\delta}}\right)^{3}}}\right).$
Appendix D Proof of Theorem 4.1
We start the proof by showing that the number of rounds where the design matrix is updated is small.
Lemma D.1 (Number of design matrix updates).
Using Algorithm 2 with a parameter $C>0$, it holds that:
$\sum_{t=1}^{T}\mathbbm{1}\{V_{t}\neq V_{t-1}\}\leq\frac{d}{\log\left({\frac{1+C}{1-1/e}}\right)}\log\left({1+\frac{T(LH_{\gamma})^{2}}{\lambda d}}\right).$
That is, the number of updates of the matrix $V_{t}$ is at most logarithmic in the number of interactions $T$.
Proof.
Denote $N_{T}=\sum_{t=1}^{T}\mathbbm{1}\left[{V_{t}\neq V_{t-1}}\right]$, and let $T^{\prime}<T$ be the last time the matrix $V_{t}$ was updated. Then,
$\det(V_{T})$ $\geq(1+C)/(1-1/e)\det(V_{T^{\prime}})$
$\geq((1+C)/(1-1/e))^{N_{T}}\det(\lambda I).$
The first inequality follows because $\det W_{T}\geq(1+C)\det V_{T^{\prime}}$ and because greedy D-Optimal design ensures that $\det V_{T}\geq\frac{1}{1-1/e}\det W_{T}$, where $W_{t}$ is defined in Algorithm 2.
Then, using the trace-determinant inequality, we have:
$\left({\frac{1+C}{1-1/e}}\right)^{N_{T}}\lambda^{d}\leq\left({\frac{\lambda d+T(LH_{\gamma})^{2}}{d}}\right)^{d}$
and then,
$N_{T}\leq\frac{d}{\log\left({\frac{1+C}{1-1/e}}\right)}\log\left({1+\frac{T(LH_{\gamma})^{2}}{\lambda d}}\right).$
We now present the proof for the regret bound, which proceeds similarly to that of Theorem 3.4 up to some modifications. The first change is in the regret decomposition, which needs to be adapted because we no longer compare to the past policy but rather to $\pi_{t}^{\prime}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}(\hat{\theta}_{t_{\operatorname{stop}}(t)},\varepsilon,\delta)$. The second change is using Lemma A.6 instead of Lemma A.5 to bound the sum of norms of trajectory features. Finally, the good events defined in the proof of Theorem 3.4 are slightly modified to account for the lazy design matrix and the new choice of comparator policy $\pi_{t}^{\prime}$.
Proof.
Let us first define the function $t_{\operatorname{stop}}:\mathbb{N}\to\mathbb{N}$, which to a time $t$, assigns the last time $t_{\operatorname{stop}}(t)\leq t$ that the update condition (line 4 in Algorithm 2) was met.
We work with the same two filtrations as before $\mathcal{F}_{t}=\sigma(x_{1},y_{1},\ldots,x_{t-1},y_{t-1})$ and $\mathcal{F}_{t}^{\theta}=\sigma(\mathcal{F}_{t},\tilde{\theta}_{t})$. And we recall that, given $\mathcal{F}_{t}$, $\tilde{\theta}_{t}$ in Algorithm 2 is sampled from $\mathcal{N}(\hat{\theta}_{t_{\operatorname{stop}}(t)},\beta_{t_{\operatorname{stop}}(t)}^{2}V_{t_{\operatorname{stop}}(t)}^{-1})$.
Step 1 (good events): Consider $\delta^{\prime}=\delta/4$, we redefine the high-probability events:
  1. 1.
    Let $\delta^{\prime\prime}=\delta^{\prime}/T$ and $c(\delta^{\prime\prime}):=\sqrt{2d\log(2d/\delta^{\prime\prime})}$. Consider the inflated ellipsoid
    $\mathcal{E}_{t_{\operatorname{stop}}(t)}^{\text{TS}}:=\left\{{\theta\in\mathbb{R}^{d}:\left\lVert\theta-\hat{\theta}_{t_{\operatorname{stop}}(t)}\right\rVert_{V_{t_{\operatorname{stop}}(t)}}\leq\beta_{t_{\operatorname{stop}}(t)}(\delta^{\prime})c(\delta^{\prime\prime})}\right\}.$
    We define the events $E_{t}:=\widehat{E}_{t}\cap\widetilde{E}_{t},\quad\widehat{E}_{t}:=\left\{{\theta^{*}\in\mathcal{E}_{t_{\operatorname{stop}}(t)}}\right\},\quad\widetilde{E}_{t}:=\left\{{\tilde{\theta}_{t}\in\mathcal{E}_{t_{\operatorname{stop}}(t)}^{\text{TS}}}\right\},$ and let $G_{1}:=\bigcap_{t=1}^{T}E_{t}$.
  2. 2.
    Let $C_{T}:=2\sqrt{\!T\!\left({\frac{8(1+C)d}{(1-1/e)^{2}}\log\left(\!1\!+\!\frac{4TL^{2}H_{\gamma}^{2}}{d\lambda}\right)\!\!+\!\frac{12dL^{2}H_{\gamma}^{2}}{\log(2)\lambda}\!\log\left(1\!+\!\frac{4L^{2}H_{\gamma}^{2}}{\log(2)\lambda}\right)}\right)}+\frac{16LH_{\gamma}}{\sqrt{\lambda}}\log\left({\frac{2}{\delta^{\prime}}}\right)$, and define $G_{2}$ as the event under which it holds that
    $\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau_{t}^{\prime})\right\rVert_{V_{t_{\operatorname{stop}}(t)}^{-1}}\mid\mathcal{F}_{t}}\right]\leq C_{T},$
    and
    $\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau_{t}^{\prime})\right\rVert_{V_{t_{\operatorname{stop}}(t)}^{-1}}\mid\mathcal{F}_{t}^{\theta}}\right]\leq C_{T}.$
  3. 3.
    Let $G_{3}=\left\{{\forall t\in[1,T]:V_{\tilde{\theta}_{t}}^{*}-V_{\tilde{\theta}_{t}}^{\pi_{t}}\leq\varepsilon}\right\}$, where $\pi_{t}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\tilde{\theta}_{t},\varepsilon,\delta^{\prime}/T}\right)$.
We also define the intersection, $G:=\bigcap_{i=1}^{3}G_{i}$, of all good events.
We now compare each of these events to their counterparts in the proof of Theorem 3.4. The event $G_{1}$ is modified because we replace $V_{t}$ by $V_{t_{\operatorname{stop}}(t)}$ and $\hat{\theta}_{t}$ by $\hat{\theta}_{t_{\operatorname{stop}}(t)}$. The event $G_{1}$ still holds with probability at least $1-2\delta^{\prime}$ using the same concentration arguments as before. The event $G_{2}$ is also modified to account for the lazy design matrix, and it holds with probability $1-\delta^{\prime}$ thanks to Lemma A.6. Finally, the event $G_{3}$ remains unchanged.
We conclude that $G$ happens with probability at least $1-\delta$.
Step 2 (regret decomposition): Since Algorithm 2 uses a different comparator policy $\pi_{t}^{\prime}$ than Algorithm 1, we derive a new regret decomposition.
$R(T)$ $=\frac{1}{2}\sum_{t=1}^{T}\left({2V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}}-V_{\theta^{*}}^{\pi^{\prime}_{t}}}\right)$
$=\frac{1}{2}\sum_{t=1}^{T}\left({(V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}})+\langle\theta^{*},\phi(\pi^{*})-\phi(\pi_{t}^{\prime})\rangle}\right)$
$=\frac{1}{2}\sum_{t=1}^{T}\left({(V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}})+\langle\theta^{*},\phi(\pi^{*})-\phi(\pi_{t})\rangle+\langle\theta^{*},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle}\right)$
$=\frac{1}{2}\sum_{t=1}^{T}\left({2(V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}})+\langle\theta^{*}-\hat{\theta}_{t_{\operatorname{stop}}(t)},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle+\langle\hat{\theta}_{t_{\operatorname{stop}}(t)},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle}\right)$
$\leq\frac{1}{2}\sum_{t=1}^{T}\left({2\underbrace{(V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}})}_{r_{t}}+\left\lVert\theta^{*}-\hat{\theta}_{t_{\operatorname{stop}}(t)}\right\rVert_{V_{t_{\operatorname{stop}}(t)}}\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t_{\operatorname{stop}}(t)}^{-1}}+\varepsilon}\right),$
where the last line follows from the Cauchy-Schwarz inequality and because $\pi_{t}^{\prime}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}(\hat{\theta}_{t_{\operatorname{stop}}(t)},\varepsilon,\delta)$.
The second term in the decomposition above can be bounded on the good event $G$ as:
$\sum_{t=1}^{T}\left\lVert\theta^{*}-\hat{\theta}_{t_{\operatorname{stop}}(t)}\right\rVert_{V_{t_{\operatorname{stop}}(t)}}\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t_{\operatorname{stop}}(t)}^{-1}}$ $\leq\beta_{T}\sum_{t=1}^{T}\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t_{\operatorname{stop}}(t)}^{-1}}$
$\leq\beta_{T}C_{T}$
where $C_{T}=2\sqrt{T\left({\frac{8(1+C)d}{(1-1/e)^{2}}\log\left(1\!+\!\frac{4TL^{2}H_{\gamma}^{2}}{d\lambda}\right)\!\!+\!\frac{12dL^{2}H_{\gamma}^{2}}{\log(2)\lambda}\log\left(1\!+\!\frac{4L^{2}H_{\gamma}^{2}}{\log(2)\lambda}\right)}\right)}+\frac{16LH_{\gamma}}{\sqrt{\lambda}}\log\left({\frac{2}{\delta^{\prime}}}\right)$.
For the first term in the regret decomposition, similarly to Appendix B, we have that:
$r_{t}=V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}}=\underbrace{\Delta_{t}(\theta^{*})-\Delta_{t}(\tilde{\theta}_{t})}_{r_{t}^{\text{TS}}}+\underbrace{\langle\tilde{\theta}_{t}-\theta^{*},\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\rangle}_{r_{t}^{\text{MLE}}}$
where we recall the gap function $\Delta_{t}(\theta):=\max_{\pi}\langle\theta,\phi(\pi)-\phi(\pi^{\prime}_{t})\rangle$.
Step 3 (bounding ${r_{t}^{\text{TS}}}$): The proof for $\sum_{t}r_{t}^{\text{TS}}$ proceeds exactly like Appendix B up to Equation (5), this is because the probability of the optimism event $O_{t}$ and the events $E_{t}$ and $\hat{E}_{t}$ is unaffected by the change to the algorithm. Then, we have that:
$R^{\text{TS}}(T)=\sum_{t=1}^{T}r_{t}^{\text{TS}}\leq\dfrac{2\beta_{T}(\delta^{\prime})c(\delta^{\prime\prime})}{p}\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t}}\right]+T\varepsilon.$
We can then conclude, on the good event $G$, that
$R^{\text{TS}}(T)\leq\dfrac{2\beta_{T}(\delta^{\prime})c(\delta^{\prime\prime})}{p}C_{T}+T\varepsilon.$
Step 4 (bounding ${r_{t}^{\text{MLE}}}$): This proof is similar to that of the first term in the regret decomposition. The only difference is that we use the confidence set for $\tilde{\theta}_{t}$ instead of $\hat{\theta}_{t_{\operatorname{stop}}(t)}$. Then, it holds that:
$\sum_{t=1}^{T}r_{t}^{\text{MLE}}$ $\leq\beta_{T}(\delta^{\prime})(1+c(\delta^{\prime\prime}))\sum_{t=1}^{T}\left\lVert\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\right\rVert_{V_{t}^{-1}}$
$\leq\beta_{T}(\delta^{\prime})(1+c(\delta^{\prime\prime}))\sum_{t=1}^{T}\mathbb{E}[\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t}^{\theta}],$
$\leq\beta_{T}(\delta^{\prime})(1+c(\delta^{\prime\prime}))C_{T}$
where the second inequality follows from the convexity of the norm.
In summary, we conclude that with probability at least $1-\delta$, the regret can be bounded as follows:
$R(T)$ $\leq\frac{1}{2}\sum_{t=1}^{T}\left({2\underbrace{(V_{\theta^{*}}^{*}-V_{\theta^{*}}^{\pi_{t}})}_{=r_{t}^{\text{TS}}+r_{t}^{\text{MLE}}}+\underbrace{\left\lVert\theta^{*}-\hat{\theta}_{t_{\operatorname{stop}}(t)}\right\rVert_{V_{t_{\operatorname{stop}}(t)}}}_{\leq\beta_{T}}\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t_{\operatorname{stop}}(t)}^{-1}}+\varepsilon}\right)$
$\leq\dfrac{2\beta_{T}(\delta^{\prime})c(\delta^{\prime\prime})}{p}C_{T}+T\varepsilon+\beta_{T}(\delta^{\prime})(1+c(\delta^{\prime\prime}))C_{T}+\frac{\beta_{T}}{2}C_{T}+T\varepsilon.$
Appendix E Discussion and background
E.1 Background on optimal experimental design
Given a possibly infinite set of features $\mathcal{X}\subset\mathbb{R}^{d}$, a D-optimal design is defined as a distribution $\pi$ such that:
$\pi\in\operatorname*{arg\,max}_{\pi\in\Delta_{\mathcal{X}}}\log\det\left({\sum_{x\in\mathcal{X}}\pi(x)xx^{\top}}\right).$
The Kiefer-Wolfowitz theorem, see [Kiefer and Wolfowitz, 1960], shows that a D-optimal design also ensures that $\max_{x\in\mathcal{X}}\left\lVert x\right\rVert_{\left({\sum_{x\in\mathcal{X}}\pi(x)xx^{\top}}\right)^{-1}}^{2}=d$. A direct consequence is that $\mathcal{X}$ is a subset of the ellipsoid $\{x\in\mathbb{R}^{d}:\left\lVert x\right\rVert_{(\sum_{x\in\mathcal{X}}\pi(x)xx^{\top})^{-1}}^{2}\leq d\}$. The Kiefer-Wolfowitz theorem can be interpreted by saying that the set $\{x\in\mathbb{R}^{d}:\left\lVert x\right\rVert_{(\sum_{x\in\mathcal{X}}\pi(x)xx^{\top})^{-1}}^{2}\leq d\}$ is the minimum volume ellipsoid containing $\mathcal{X}$, see [Lattimore and Szepesvári, 2020, Theorem 21.1].
In the case where we have a budget of $n$ samples for the design, we can define D-optimal design as the best allocation $\{n_{x}\}_{x\in\mathcal{X}}\in\mathbb{N}$ of the $n$ samples such that $\log\det\left({\sum_{x\in\mathcal{X}}n_{x}xx^{\top}}\right)$ is maximized and $\sum_{x\in\mathcal{X}}n_{x}=n$. Finding a D-optimal design with a budget of $n$ is a challenging combinatorial optimization problem. However, there is a key property of the $\log\det$ function that enables an efficient approximation scheme. Namely, for a set $\mathcal{X}\subset\mathbb{R}^{d}$, the set function $S\subset\mathcal{X}\to\log\det\left({\lambda I+\sum_{x\in S}xx^{\top}}\right)$ is submodular if $\lambda$ is greater than one. Submodularity is a property describing decreasing additional benefit, known as diminishing returns. Fortunately, a submodular set function can be approximately maximized using a greedy algorithm, [Nemhauser et al., 1978]. Therefore, using Algorithm 3, we know that if $t_{\operatorname{stop}}$ and $t_{\operatorname{stop}}^{\prime}$ are two consecutive update times, then the design matrix satisfies: $\log\det\left({V_{t_{\operatorname{stop}}^{\prime}}}\right)\geq(1-1/e)\max_{S\subset\mathcal{X}}\log\det\left({V_{t_{\operatorname{stop}}}+\sum_{x\in S}xx^{\top}}\right)$.
E.2 On the intractability of optimistic approaches
Optimism in the face of uncertainty is a widely used principle for regret minimization and exploration in reinforcement learning. In the bandit setting, optimistic algorithms can be applied directly and yield minimax-optimal regret bounds [Auer, 2002]. However, in more general settings–such as linear bandits or MDPs–these approaches often lead to intractable optimization problems.
Optimism for regret minimization. For regret minimization in RLHF, an optimistic algorithm would choose the policy $\pi_{t}$ to maximize the upper confidence bound on the reward difference relative to a comparator policy $\pi_{t}^{\prime}$, typically the previous policy:
$\pi_{t}=\operatorname*{arg\,max}_{\pi}\max_{\theta\in\mathcal{E}_{t}}V_{\theta}^{\pi}-V_{\theta}^{\pi_{t}^{\prime}}=\operatorname*{arg\,max}_{\pi}V_{\hat{\theta}_{t}}^{\pi}+\beta_{t}\left\lVert\phi(\pi)-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}},$
where $\mathcal{E}_{t}$ is a confidence set for $\theta^{*}$. This leads to the following bound on the instantaneous regret:
$r_{t}$ $=\left({V^{*}_{\theta^{*}}-V^{\pi_{t}^{\prime}}_{\theta^{*}}}\right)-\left({V^{\pi_{t}}_{\theta^{*}}-V^{\pi_{t}^{\prime}}_{\theta^{*}}}\right)$
$\leq\left({\max_{\pi}\max_{\theta\in\mathcal{E}_{t}}V_{\theta}^{\pi}-V_{\theta}^{\pi_{t}^{\prime}}}\right)-\left({V^{\pi_{t}}_{\theta^{*}}-V^{\pi_{t}^{\prime}}_{\theta^{*}}}\right)$
$=\left({V_{\tilde{\theta}_{t}}^{\pi_{t}}-V_{\tilde{\theta}_{t}}^{\pi_{t}^{\prime}}}\right)-\left({V^{\pi_{t}}_{\theta^{*}}-V^{\pi_{t}^{\prime}}_{\theta^{*}}}\right)$
$\leq\left\lVert\tilde{\theta}_{t}-\theta^{*}\right\rVert_{V_{t}}\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\leq 2\beta_{t}(\delta^{\prime})\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}},$
where $\tilde{\theta}_{t}\in\mathcal{E}_{t}$. The cumulative regret can then be bounded using standard elliptical potential arguments (see Lemma A.5).
Optimism for pure exploration. Optimistic algorithms are also successful for minimizing the suboptimality gap of the final policy. In this case, algorithms aim to maximize the information gain by selecting either a policy pair $(\pi_{t},\pi_{t}^{\prime})$ or a single policy $\pi_{t}$ relative to a reference policy $\pi_{t}^{\prime}$ to maximize the exploration bonus:
$\left\lVert\phi(\pi)-\phi(\pi^{\prime})\right\rVert_{V_{t}^{-1}}.$
These formulations are appealing due to their theoretical guarantees. However, not much can be said about the computational efficiency of the subsequent optimization problems.
Challenge of optimizing the exploration bonus. As shown by Efroni et al. [2021], optimizing these exploration bonuses is intractable even in tabular MDPs with a known transition law. Specifically, maximizing the bonus reduces to the following problem over the set of discounted occupancy measures:
$\max_{\mu\in\mathcal{M}}\langle r_{\theta},\mu\rangle+\beta_{t}\left\lVert\mathbb{E}_{(s,a)\sim\mu}\phi(s,a)-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}},$
where :=$\mathcal{M}:=\left\{{\mu^{\pi}:\mu^{\pi}(s,a)=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{P}_{\pi}(s_{t}=s,a_{t}=a),\pi:\mathcal{S}\to\Delta_{\mathcal{A}}}\right\}$ is the set of valid occupancy measures induced by stationary policies [Puterman, 2014]. The norm term is strictly convex in $\mu$ due to the strict convexity of the norm and the linearity of expectations. As a result, this is a convex maximization problem, which is NP-hard in general [Atamtürk and Gómez, 2017].
These intractability results highlight a major limitation of optimistic algorithms. Although they yield strong statistical guarantees, their practical implementation in general MDPs remains computationally challenging. This motivates exploring alternative approaches, such as randomized exploration, which can provide almost optimal theoretical guarantees while being computationally tractable.
Appendix F Experiments
Here, we provide additional details for experiments on Isaac Lab’s Isaac-Cartpole-v0 environment, as well as experimental results for the pure exploration version of our algorithm.
Implementation details
All experiments run on Isaac Lab’s unmodified Isaac-Cartpole-v0 environment using the default PPO configuration. We train over 30 RLHF iterations, using 30 steps of PPO at each iteration, and training is repeated for 10 independent seeds. For the randomized exploration, we set $\beta_{t}=0.01+\max(1,\log t)$ and $\lambda=1$, and for lazy updates we set $C=1$. At each RLHF iteration we compare 5 independently sampled trajectories of horizon $H=150$. For the maximum likelihood estimation we perform 500 Adam steps (batch size $64$, $\ell_{2}$ penalty $\lambda=10^{-1}$). Experiments were executed on a single machine equipped with an Intel i9-14900KS CPU and an NVIDIA RTX 4090 GPU; completing 30 RLHF iterations required approximately 2 min 50 s.
Pure exploration
Our results for the pure exploration setting with $\alpha=0$ are shown in Figure 2 and 3 below. In Figure 2, we see that all three versions of RPO achieve performance competitive to RL with the ground truth reward [^4] Note that the RL baseline makes $20024\times 4096\times 50=4^{\prime}915^{\prime}$ 200 queries to the ground truth reward, whereas RPO uses at most 150 binary preference queries., but LRPO-OD needs the least preference queries. Moreover, in Figure 3 we see that the performance during training is poor, i.e. the regret is large, which is to be expected due to the pure exploration scheme.
Figure 2: Comparison of RLHF algorithms in terms of (a) the last iterate ground truth reward $V_{\theta^{*}}^{\hat{\pi}}$ (estimated from samples) and (b) number of preference queries performed. In particular, for the choice $\alpha=0$, we compare RPO (green, Algorithm 1) with it’s lazy versions LRPO and LRPO-OD (orange & blue). Here, LRPO and LRPO-OD refer to Algorithm 2 without and with optimal design subroutine. The error bars indicate the 0.2 and 0.8 quantiles, across 10 independent runs. The dashed blue line indicates the mean reward achieved by PPO with the ground truth parameter $\theta^{*}$.
Figure 3: Comparing the ground truth rewards $V_{\theta^{*}}^{\pi_{t}}$ (estimated from samples) of RLHF algorithms for pure exploration (i.e. $\alpha=0$) during training, using the same color codes as in Figure 2.